### 19 Mar 2026 v2

# On K-peak solutions for the Yamabe equation on product manifolds

Abstract.
Let $(M^{n},g)$ and $(X^{m},h)$ be closed manifolds $m,n>2$, such that $(X,h)$ has constant positive scalar curvature. We consider the one parameter family of products $(M\times X,g+\epsilon^{2}h)$, $\epsilon>0$. Let $\xi_{0}\in M$ be a stable, isolated, critical point of a function $\Phi:M\rightarrow{\mathbb{R}}$, eq. (9), that depends on the scalar curvature, the dimensions $m,n$, the norm of the Ricci curvature and the norm of the curvature tensor of $g$. We prove that, if either the scalar curvature of $g$ is constant or a certain dimensional constant $\beta=0$, then for each $K\in{\mathbb{N}}$, there is some $\epsilon_{0}>0$ such that for every $\epsilon\in(0,\epsilon_{0})$ the subcritical Yamabe equation $\epsilon^{2}\Delta u+(1+{\bf{c}}\epsilon^{2}s_{g})u=u^{q}$ on $(M\times X,g+\epsilon^{2}h)$ has a positive $K$-peak solution, which concentrates around $\xi_{0}$. Where, ${\bf{c}}=\frac{N-2}{4(N-1)}$, $N=n+m$, $q=\frac{N+2}{N-2}$ and $s_{g}$ the scalar curvature of $(M,g)$. This covers some remaining cases of previous results [19] and gives examples of multiplicity of positive solutions for the Yamabe equation.
Key words and phrases:
Yamabe equation
1991 Mathematics Subject Classification:
35J20, 35J60, 53C21
ENES UNAM. Departamento de Matemáticas. León, Gto., México. mruiz@enes.unam.mx
ENES UNAM. Departamento de Matemáticas. León, Gto., México. areli@enes.unam.mx.
* Corresponding author.
1. Introduction
Let $(M^{n},g)$ be a closed manifold $n>2$. The Yamabe problem consists of finding a metric of constant scalar curvature in the conformal class of the metric $g$. If one writes a conformal metric to $g$ as $\tilde{g}=u^{\frac{4}{n-2}}g$, where $u\in C^{\infty}(M)$ and $u>0$, the resulting metric $\tilde{g}$ will have constant scalar curvature $s_{\tilde{g}}=\lambda$ if and only if $u$ solves the Yamabe equation:
(1) $a_{n}\Delta u+s_{g}u=\lambda u^{p_{n}-1}$
where $a_{n}=\frac{4(n-1)}{n-2}$ and $p_{n}=\frac{2n}{n-2}$. It is a classical result, solved in various steps by H. Yamabe [23], N.S. Trudinger [22], T. Aubin [2], and R. Schoen [21], that such a postive, smooth, function $u$ always exists (see [16], [1], for a detailed discussion of the Yamabe problem). It has been of interest to ask whether other metrics of constant scalar curvature exist in the same conformal class, this is, whether problem (1) has a unique solution. Uniqueness in this sense is known to be true for some special cases, for example, when the original metric has non-positive scalar curvature, by the maximum principle [2], or when the original metric is an Einstein metric, not isometric to the sphere with the round metric, by the work of M. Obata [13]. On the other hand, when the original metric has positive scalar curvature, many examples of different metrics with constant scalar curvature and unit volume have been found in the same conformal class, see for instance [3], [6], [10], [14], [15], [17], [18], [20].
Let $(M^{n},g)$ and $(X^{m},h)$ be closed manifolds $m,n>2$, such that $(X,h)$ has constant positive scalar curvature. Consider the one parameter family of products $(M\times X,g+\epsilon^{2}h)$. The Yamabe equation on this family is
(2) $a_{N}(\Delta_{g}+\Delta_{\epsilon^{2}h})u+(s_{g}+\epsilon^{-2}s_{h})u=u^{p_{N}-1}$
where $N=m+n$. One may renormalize equation (2) so that $s_{h}=a_{N}$ and look only for solutions that depend solely on $M$, $u:M\rightarrow{\mathbb{R}}$. In this case, equation(2) is equivalent to
(3) $\Delta_{g}u+({\bf{c}}s_{g}\epsilon^{-2}+1)u=u^{p_{N}-1}$
where ${\bf{c}}=c_{N}=a_{N}^{-1}$. To solve (3), we follow the techniques of the Lyapunov-Schmidt reduction, which have been applied by many authors for these type of equations. We refer the reader, in particular, to the work of Dancer, Micheletti and Pistoia [4], of Rey and Ruiz [19] and of Deng, Khemiri and Mahmoudi [7].
We model the solutions to (3) after the solutions of a limit equation in ${\mathbb{R}}^{n}$,
(4) $\Delta U+U=U^{p-1}$
where $p=p_{N}$, $2<p<\frac{2n}{n-2}$ and $n>2$. It is well known, by the work of Kwong, [11], that there exists a unique, positive, spherically symmetric, function $U\in H^{1}({\mathbb{R}}^{n})$ such that $U$ is a solution to (4). Moreover, $U$ and its derivatives are exponentially decaying at infinity, this is
(5) $\lim_{|x|\rightarrow\infty}U(|x|)|x|^{\frac{n-1}{2}}e^{|x|}=c>0,\ \ \ \lim_{|x|\rightarrow\infty}U^{\prime}(|x|)|x|^{\frac{n-1}{2}}e^{|x|}=-c.$
Through the article we will denote this solution by $U$, assuming $n$ and $p$ are clear from the context. Using this model solution in ${\mathbb{R}}^{n}$ we will construct a multipeak soultion on the manifold in the following way.
For any $\epsilon>0$, note that $U_{\epsilon}=U(\frac{x}{\epsilon})$ is a solution to
(6) $\epsilon\Delta U_{\epsilon}+U_{\epsilon}=U_{\epsilon}^{p-1}.$
Now, since $M$ is closed, we can fix $r_{0}>0$ such that for any $x\in M$, the exponential map $exp_{x}:T_{x}M\rightarrow M$, at $x$ restricted to $B(0,r_{0})$, is a diffeomorphism. $B(0,r_{0})$ denotes a ball in ${\mathbb{R}}^{n}$ centered at 0, with radius $r_{0}$ and we will denote by $B_{g}(x,r)$ the geodesic ball in $M$ centered at $x$ with radius $r$.
For fixed $r<r_{0}$, we then define on $M$ the function $W_{\epsilon,\xi}:M\rightarrow{\mathbb{R}}$. Let $\chi_{r}$ be a smooth, radial cutoff function in ${\mathbb{R}}^{n}$, $\chi_{r}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ such that $\chi(z)=1$ if $z\in B(0,\frac{r}{2})$ and $\chi(z)=0$ if $z\in{\mathbb{R}}^{n}\setminus B(0,r)$. Let
(7) $W_{\epsilon,\xi}(x)=\begin{cases}U_{\epsilon}(exp_{\xi}^{-1}(x))\chi_{r}(exp_{\xi}^{-1}(x))&\textit{if}\ \ x\in B_{g}(\xi,r)\\ 0&\textit{otherwise}\end{cases}$
Note that $W_{\epsilon,\xi}(x)$ is an approximate solution to (3), which concentrates around $\xi$ as $\epsilon\rightarrow 0$, by the exponential decay of $U$, eventually getting closer to an exact solution. Adding $K$ of these peaks $W_{\epsilon,\xi}(x)$, centered at different $K$ strategic points $\xi_{1},\xi_{2},\dots,\xi_{K}\in M$, we get an approximate solution of (3) which has $K$-peaks: $\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}(x)$. Then, by adding some perturbations of these approximate solutions we will get an exact solution to equation (3).
For $\mathbf{c}=c_{N}$, we define
(8) $\beta:=\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}(z)-\frac{1}{n(n+2)}\int_{{\mathbb{R}}^{n}}|\nabla U(z)|^{2}|z|^{2}dz$
note that $\beta$ is a dimensional constant that depends only on $n$ and $m$. The precise values of $\beta$ for general $m,n$ are not known, to the best of our knowledge. Analytical computations have been done for example for the case $m+n=4$, in [5]. On the other hand, in [19] the authors provided numerical computations for values of $m$ and $n$ such that $m+n\leq 9$. In all these cases $\beta<0$.
If $\beta\neq 0$ and $s_{g}$ has isolated, stable, critical points, multipeak solutions like the ones described above were studied in [19]. In this case, the solutions concentrate around isolated critical points of the scalar curvature $s_{g}$. In the present article we work on some of the remaining cases. In particular, we show that, if either $\beta=0$ or $s_{g}$ is constant, $K$-peak solutions concentrate around an isolated, stable, critical point of the functional $\Phi:M\rightarrow{\mathbb{R}}^{n}$, given by
[…]
where $s_{g}$, $Ric_{\xi}$ and $R_{\xi}$ are the scalar curvature, Ricci curvature and tensor of curvature of $(M,g)$, respectively. Meanwhile $c_{1},c_{6},c_{7},c_{8}$ and $c_{9}$ are constants that depend on the dimensions $m,n$, and are given in equations (69) through (74). Note that, for example, if $(M,g)$ is an Einstein manifold, then critical points of $R_{\xi}||^{2}$ would be critical points of $\Phi(\xi)$. We also remark that at least the constant $c_{1}=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz\neq 0$ for any values of $n,m>2$.
Theorem 1.1.
Suppose that $\beta=0$ or that $s_{g}$ is constant. Let $\xi_{0}$ be an isolated local $C^{1}$ stable critical point of the functional $\Phi(\xi)$. Then, for each positive integer $K$, there exists $\epsilon_{0}=\epsilon_{0}(K)>0$ such that for each $\epsilon\in(0,\epsilon_{0})$ there are points $\xi_{1}^{\epsilon},\xi_{2}^{\epsilon},\dots,\xi_{K}^{\epsilon}\in M$, so that
$\frac{d_{g}(\xi_{i}^{\epsilon},\xi_{j}^{\epsilon})}{\epsilon}\rightarrow\infty\textit{\ and \ }d_{g}(\xi_{0}^{\epsilon},\xi_{j}^{\epsilon})\rightarrow 0$
and a solution $u_{\epsilon}$ of problem (3), so that
$u_{\epsilon}-\sum_{i=1}^{K}W_{\epsilon,\xi_{j}^{\epsilon}}||_{\epsilon}\rightarrow 0$
Our approach follows some ideas in [8], like expanding the energy functional to order 4, which gives place to the functional $\Phi(\xi)$. We then use Lyapunov-Schmidt reduction techinques for subcritical elliptic PDEs on manifolds, like the ones applied, for example, in [12], [7]; in particular [4] for multipeak solutions and [19] for the subcritical Yamabe equation.
In Section 2 we introduce some notation and background, as well as a way to improve the aproximate solution to higher order. In section 3 we prove Theorem 1.1, assuming propositions 3.1, 3.2 and 3.3. Section 4 is dedicated to compute the expansion of the energy functional, which is the content of proposition 3.2. In section 5 we prove proposition 3.1, regarding the finite dimensional reduction. Finally, in section 6, we provide some technical estimates and the proof of proposition 3.3.
Acknowledgments. The authors were supported by grant UNAM-DGAPA-PAPIIT IN108824 and grant UNAM-DGAPA-PAPIIT IN108224, respectively.
2. Approximate solutions
Critical points of the functional $E:H^{1}({\mathbb{R}}^{n})\rightarrow{\mathbb{R}}$
$E(f)=\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}|\nabla f|^{2}+\frac{1}{2}f^{2}-\frac{1}{p}(f^{+})^{p}\right)dx$
are positive solutions to the limit equation (4). We denote by $S_{0}=\nabla E:H^{1}({\mathbb{R}}^{n})\rightarrow H^{1}({\mathbb{R}}^{n})$. Note that $S_{0}(U)=0$ and that $U$ is non degenerate in the sense that $Kernel(S_{0}^{\prime}(U))$ is spanned by
$\psi^{i}(x):=\frac{\partial U}{\partial x_{i}}(x)$
for $i=1,2,\dots,n$. This implies that the solution $U_{\epsilon}$ to equation (6) is a critical point of the functional
$E_{\epsilon}(F)=\epsilon^{-n}\int_{{\mathbb{R}}^{n}}\left(\frac{\epsilon^{2}}{2}|\nabla f|^{2}+\frac{1}{2}f^{2}-\frac{1}{p}(f^{+})^{p}\right)dx$
We denote by ${S_{0}}_{\epsilon}=\nabla E_{\epsilon}$. $U_{\epsilon}$ is non degenerate in the sense that $Kernel({S_{0}}_{\epsilon}^{\prime}(U_{\epsilon}))$ is spanned by
$\psi_{\epsilon}^{i}(x):=\psi^{i}\left(\frac{x}{\epsilon}\right)$
for $i=1,2,\dots,n$. On $M$, for $i=1,2,\dots,n$, we then may define
(10) $Z^{i}_{\epsilon,\xi}(x):=\begin{cases}\psi^{i}_{\epsilon}(exp_{\xi}^{-1}(x))\chi_{r}(exp_{\xi}^{-1}(x))&\textit{if}\ \ x\in B_{g}(\xi,r)\\ 0&\textit{otherwise}\end{cases}$
Let $H_{\epsilon}$ be the Hilbert space $H_{g}^{1}(M)$ with the inner product
$\langle u,v\rangle_{\epsilon}:=\frac{1}{\epsilon^{n}}\left(\epsilon^{2}\int_{M}\nabla_{g}u\nabla_{g}v\ d\mu_{g}+\int_{M}(1+\epsilon^{2}\mathbf{c}s_{g})uv\ d\mu_{g}\right)$
which induces the norm
$u||^{2}_{\epsilon}:=\frac{1}{\epsilon^{n}}\left(\epsilon^{2}\int_{M}|\nabla u|^{2}d\mu_{g}+\int_{M}(1+\epsilon^{2}\mathbf{c}s_{g})u^{2}d\mu_{g}\right)$
Recall ${\bf c}=c_{N}=a_{N}^{-1}=\frac{N-2}{4(N-1)}$, $N=n+m$. Note that for any function $f$, with $f_{\epsilon}(x)=f(\frac{x}{\epsilon})$, $f_{\epsilon}||_{\epsilon}$ is independent of $\epsilon$.
For $\epsilon>0$ and $\bar{\xi}=(\xi_{1},\xi_{2},\dots\xi_{K})\in M^{K}$, let
$K_{\epsilon,\bar{\xi}}:=\text{span}\left\{Z^{i}_{\epsilon,\xi_{j}}|\ i=1,\dots,n\ \text{and}\ j=1,\dots,K\right\}$
$K_{\epsilon,\bar{\xi}}^{\perp}:=\left\{\phi\in H^{1}_{\epsilon}|\ \langle\phi,Z^{i}_{\epsilon,\xi_{j}}\rangle_{\epsilon},i=1,\dots,n\ \text{and}\ j=1,\dots,K\right\}$
Let $\Pi_{\epsilon,\xi}:H_{\epsilon}\rightarrow K_{\epsilon,\bar{\xi}}$ and $\Pi_{\epsilon,\xi}^{\perp}:H_{\epsilon}\rightarrow K_{\epsilon,\bar{\xi}}^{\perp}$ be orthogonal projections.
Let $L_{\epsilon}^{p}$ denote the Banach space $L_{g}^{p}(M)$ with the norm $|u|_{p,\epsilon}:=\left(\frac{1}{\epsilon^{n}}\int_{M}|u|^{p}d\mu_{g}\right)^{1/p}$. Since $2<p<\frac{2n}{n-2}$, it follows from the usual Sobolev embeddings, that there is a constant $c$ independent of $\epsilon$ such that
[…]
In particular, the embedding $H_{\epsilon}\xhookrightarrow{}L_{\epsilon}^{p}$ is a compact continuous map. We denote $p^{\prime}:=\frac{p}{p-1}$. We identify the dual space $L^{p}_{\epsilon}$ with $L^{p^{\prime}}_{\epsilon}$, with the pairing
$\langle\varphi,\psi\rangle=\frac{1}{\epsilon^{n}}\int_{M}\varphi\psi$
for $\varphi\in L^{p}_{\epsilon}$, $\psi\in L^{p^{\prime}}_{\epsilon}$. The adjoint operator $i^{*}_{\epsilon}:L_{\epsilon}^{p^{\prime}}\rightarrow H_{\epsilon}$ is a continuous map so that
(12) $u=i^{*}_{\epsilon}(v)\iff\langle i^{*}_{\epsilon}(v),\varphi\rangle_{\epsilon}=\frac{1}{\epsilon^{n}}\int_{M}v\varphi,\varphi\in H_{\epsilon}\iff-\epsilon^{2}\Delta_{g}u+(1+{\bf c}s_{g}\epsilon^{2})u=v\text{\ (weakly) on\ }M\\$
For the same constant as in (11), for any $v\in L^{p^{\prime}}_{\epsilon}$,
(13) $i^{*}_{\epsilon}(v)||_{\epsilon}\leq c|v|_{p^{\prime},\epsilon}$
Let $f(u):=(u^{+})^{p-1}$. Note that
(14) $S_{\epsilon}(u)=u-i^{*}_{\epsilon}(f(u)),u\in H_{\epsilon}$
so that problem (3) can be rewritten as
(15) $u=i^{*}_{\epsilon}(f(u)),u\in H_{\epsilon}$
The following will be useful. They are computed, for example, in [9].
Lemma 2.1.
The following expansions hold.
  1. i)
    In a normal coordinates neighborhood of $\xi\in M$,
    (16) $\Delta_{g}u=\Delta_{{\mathbb{R}}^{n}}u-\frac{\epsilon^{2}}{3}R_{kijl}z_{k}z_{l}\partial_{ij}^{2}u+\frac{2\epsilon^{2}}{3}R_{kssj}z_{k}\partial_{j}u+o(\epsilon^{2}|z|^{2})$
    (17) $g_{ij}=\delta_{ij}-\frac{\epsilon^{2}}{3}R_{kijl}z_{k}z_{l}-\frac{\epsilon^{3}}{6}\nabla_{m}R_{kijl}z_{k}z_{l}z_{m}-\epsilon^{4}\left(\frac{1}{20}\nabla_{pq}R_{kijl}+\frac{2}{45}R_{kilr}R_{pjqr}\right)z_{k}z_{l}z_{p}z_{q}$
    +ϵ49R_kislR_psjqz_kz_lz_pz_q+o(ϵ^4—z—^4)
    (18) $\sqrt{det(g)}=1-\epsilon^{2}\frac{1}{6}R_{kl}z_{k}z_{l}-\frac{\epsilon^{3}}{12}\nabla_{m}R_{kl}z_{k}z_{l}z_{m}-\epsilon^{4}\left(\frac{1}{40}\nabla_{pq}R_{kl}+\frac{1}{180}R_{kilr}R_{piqr}-\frac{1}{72}R_{kl}R_{pq}\right)z_{k}z_{l}z_{p}z_{q}$
    +o(ϵ^4—z—^4)
  2. ii)
    Let $w(z)=u(\exp_{\xi}(z))$, $z\in B(0,r)\subset{\mathbb{R}}^{n}$, $\xi\in M$,
    (19) $\Delta_{{\mathbb{R}}^{n}}w=\Delta_{g_{\xi}}w+(g^{ij}_{\xi}-\delta_{ij})\partial^{2}_{ij}w-\epsilon^{2}g^{ij}_{\xi_{i}}\Gamma_{ij}^{k}\partial_{k}w+o(\epsilon^{2})$
    where $R_{ijkl}$ and $R_{kl}$ denote the curvature tensor and the Ricci curvature.
We will use a second order improvement of the approximation solution $W_{\epsilon,\xi}$ of the form
(20) ${u_{1}}_{\epsilon,\xi}=W_{\epsilon,\xi}+\epsilon^{2}V_{\epsilon,\xi}$
where $V_{\epsilon,\xi}$ will be a function $V_{\epsilon,\xi}:M\rightarrow{\mathbb{R}}$, such that the second order terms of $\epsilon$ vanish for
$\Delta_{g}{u_{1}}_{\epsilon,\xi}+(1+\epsilon^{2}{\bf c}s_{g}){u_{1}}_{\epsilon,\xi}-|{u_{1}}_{\epsilon,\xi}|^{p-1}$
More precisely, let
(21) $V_{\epsilon,\xi}(x):=\begin{cases}V_{\epsilon}(exp_{\xi}^{-1}(x))\chi_{r}(exp_{\xi}^{-1}(x))&\textit{if}\ \ x\in B_{g}(\xi,r)\\ 0&\textit{otherwise}\end{cases}$
where $V_{\epsilon}(z):=V(\frac{z}{\epsilon})$ and $V:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is the unique bounded solution of the problem
(22) $L_{0}v=-\frac{1}{3}R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U-{\bf c}s_{g}U$
where $L_{0}$ is the linear operator defined by,
(23) $L_{0}v:=-\Delta_{{\mathbb{R}}^{n}}v+v-(p-1)|U|^{p-2}v$
Since $U$ is radial,
$\partial_{i}U=z_{i}\frac{U^{\prime}}{|z|}\textnormal{\ \ \ and \ \ \ }\partial_{ij}^{2}U=\delta_{ij}\frac{U^{\prime\prime}}{|z|^{2}}+z_{i}z_{j}\frac{U^{\prime}}{|z|}-z_{i}z_{j}\frac{U^{\prime}}{|z|^{3}}$
then
$R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U=R_{kiil}z_{k}z_{l}\frac{U^{\prime}}{|z|}+R_{kijl}z_{k}z_{l}z_{i}z_{j}\frac{U^{\prime\prime}}{|z|^{2}}-R_{kijl}z_{k}z_{l}z_{i}z_{j}\frac{U^{\prime}}{|z|^{3}}$
with the second and third terms being zero, by the antisymmetry of the curvature tensor. Thus, we have,
$\frac{1}{3}R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U=\frac{1}{3}R_{kl}z_{k}\ \partial_{l}U$
this implies that eq. (22) is equivalent to
(24) $L_{0}v=\frac{1}{3}R_{kl}z_{k}\ \partial_{l}U-{\bf c}s_{g}U$
Since $L_{0}$ is linear, one can look for a solution $V=V_{1}+V_{2}$ such that
(25) $L_{0}V_{1}=\frac{1}{3}R_{kl}\ z_{k}\ \partial_{l}U$
and
(26) $L_{0}V_{2}=-{\bf c}s_{g}U$
Note that (25) is always solvable since the right-hand side is orthogonal to the Kernel of $L_{0}$ (which is spanned by $\partial_{i}U$, $i=1,2,\dots,n$). Moreover, since $U$ is radial, then $\partial_{l}U=\frac{U^{\prime}}{|z|}z_{l}$, hence eq. (25) is equivalent to
(27) $L_{0}V_{1}=\frac{1}{3}R_{kl}\frac{U^{\prime}}{|z|}z_{k}z_{l}$
By direct computation, for a radial function $\varphi=\varphi(|z|)$ on ${\mathbb{R}}^{n}$,
$L(\varphi z_{k}z_{l})=\left(\varphi^{\prime\prime}+(N+3)\frac{\varphi^{\prime}}{|z|}+\varphi-(p-1)U^{p-2}\varphi\right)z_{k}z_{l}$
Then, one can look for a radial function $\psi=\psi(|z|)$, such that
(28) $L(\psi z_{k}z_{l})=\frac{U^{\prime}}{|z|}z_{k}z_{l}$
since $\frac{U^{\prime}}{|z|}$ is radial. In this way, the solution to (27) will be given by
(29) $V_{1}=\frac{1}{3}R_{kl}\ \psi\ z_{k}z_{l}$
Moreover, by the form of the operator $L_{0}$ (eq. (23)) and of the right hand side of (27), $V_{1}$ and its derivatives are exponentially decaying at infinity.
On the other hand, by direct computation
(30) $L_{0}(U^{\prime}|z|)=-2\Delta U\textnormal{\ \ \ and \ \ \ }L_{0}(U)=(2-p)U^{p-1}$
Hence,
(31) $V_{2}={\bf c}s_{g}\left(\frac{1}{2}U^{\prime}|z|-\frac{1}{2-p}U\right)$
satisfies (26). Of course, this implies that $V=V_{1}+V_{2}$ satisfies (24) and, moreover, that $V$ and its derivatives are exponentially decaying at infinity.
For $\xi\in M$ and $\epsilon>0$, let
(32) $Y_{\epsilon,\xi}:=W_{\epsilon,\xi}+\epsilon^{2}V_{\epsilon,\xi}$
and, for $\xi_{1},\xi_{2},\dots,\xi_{K}\in M$, with $\bar{\xi}=(\xi_{1},\xi_{2},\dots,\xi_{K})$,
(33) $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)$
Let $\xi_{0}\in M$ an isolated, stable, $C^{1}$ critical point of the functional $\Phi$ (eq. (9)). Let $K\geq 1$ be an integer. Given $\rho>0$, $\epsilon>0$, let
(34) $D_{\epsilon,\rho}^{K}:=\{\bar{\xi}\in M^{K}|\ d_{g}(\xi_{0},\xi_{i})<\rho,\ i=1,2,\dots,K,\ \sum_{i\neq j}^{K}\left(U_{\epsilon}(\exp_{\xi_{i}}^{-1}\xi_{j})\right)<\epsilon^{4}\}$
Note that $\bar{\xi}_{\epsilon}\in D_{\epsilon,\rho}^{K}$ implies, for $i\neq j$,
(35) $\lim_{\epsilon\rightarrow 0}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}=+\infty$
This follows from the fact that
$U_{\epsilon}(\exp_{\xi_{i}}^{-1}\xi_{j})=U\left(\frac{\exp_{\xi_{i}}^{-1}\xi_{j}}{\epsilon}\right)=U\left(\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)<\epsilon^{4}$
and $U$ is a radial, positive, exponentially decreasing function. Moreover, it follows from (5) that, for any $\delta>0$ we have
(36) $\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon^{4}}e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}=0$
since, otherwise, if for some $a>0$, $\lim_{\epsilon\rightarrow 0}e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4}$. Then, for some small and positive values of $\epsilon$ ($\epsilon\rightarrow 0$),
$\lim_{\epsilon\rightarrow 0}e^{-\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4}e^{\delta\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}.$
From this, applying (5) with $|x|=\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}$, and since $U\left(\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)<\epsilon^{4}$:
(37) $\epsilon^{4}\left(\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)^{\frac{n-1}{2}}>ce^{-\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>ca\epsilon^{4}e^{\delta\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}},$
which gives a contradiction as $\epsilon\rightarrow 0$, since one side has polynomial growth and the other one exponential growth. In a similar way, note that, for $\delta>0$ small, also
(38) $\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon^{4}}e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}=0$
since, otherwise, if for some $b>0$, $\lim_{\epsilon\rightarrow 0}e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4}$, for some small and positive values of $\epsilon$, as $\epsilon\rightarrow 0$. And then, for those values,
(39) $\lim_{\epsilon\rightarrow 0}e^{-\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4}e^{(\frac{1}{2}-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}.$
From this, applying (5) with $|x|=\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}$, and since $U$ is decreasing,
$U\left(\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)\leq U\left(\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)<\epsilon^{4},$
hence
(40) $\epsilon^{4}\left(\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)^{\frac{n-1}{2}}>ce^{-\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>cb\epsilon^{4}e^{(1-2\delta)\left(\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}\right)},$
which gives a contradiction, for small $\delta$, as $\epsilon\rightarrow 0$, since one side has polynomial growth and the other, exponential growth. Equations (36) and (38) imply that $e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ and $e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ are $o(\epsilon^{4})$ for ${\xi_{\epsilon}}_{i}$, ${\xi_{\epsilon}}_{j}\in\bar{\xi}_{\epsilon}$, $i\neq j$, with $\bar{\xi}_{\epsilon}\in D^{k}_{\epsilon,\rho}$.
We will look for a solution to (3) of the form
(41) $u_{\epsilon}=Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}$
where $Y_{\epsilon,\bar{\xi}}$ is defined in (33), $\bar{\xi}_{\epsilon}\in D^{k}_{\epsilon,\rho}$ and the rest term $\phi_{\epsilon,\bar{\xi}}\in K^{\perp}_{\epsilon,\bar{\xi}}$.
Positive solutions to problem (3) are critical points of the functional $J_{\epsilon}:H^{1}(M)\rightarrow{\mathbb{R}}$, given by
(42) $J_{\epsilon}(u)=\frac{1}{\epsilon^{n}}\int_{M}\left(\frac{\epsilon^{2}}{2}|\nabla u|^{2}+\frac{1}{2}(1+{\bf c}s_{g}\epsilon^{2})u^{2}-\frac{1}{p_{N}}(u^{+})^{p_{N}}\right)d\mu_{g}$
with $u^{+}(x)=\max\{u(x),0\}$.
Let $S_{\epsilon}=\nabla J_{\epsilon}:H_{\epsilon}\rightarrow H_{\epsilon}$. To solve problem (3) we must solve $S_{\epsilon}(u_{\epsilon})=0$. Or in an equivalent way, we must solve the system,
(43) $\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})\right\}=0.$
(44) $\Pi_{\epsilon,\xi}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})\right\}=0.$
Note that (44) is a finite dimensional problem.
3. Existence result (proof of Theorem 1.1)
We will use the following to prove Theorem 1.1. The proof of this proposition is postponed to section 5.
Proposition 3.1.
There exist $\rho_{0}>0$, $\epsilon_{0}>0$, $c>0$ and $\sigma>0$ such that for any $\rho\in(0,\rho_{0})$, $\epsilon\in(0,\epsilon_{0})$ and $\bar{\xi}\in D_{\epsilon,\rho}^{K}$ there is a unique $\phi_{\epsilon,\bar{\xi}}=\phi(\epsilon,\bar{\xi})\in K_{\epsilon,\bar{\xi}}^{\perp}$, which solves equation (43) and satisfies
(45) $\phi_{\epsilon,\bar{\xi}}||_{\epsilon}\leq c\left(\epsilon^{3}+\sum_{i\neq j}e^{-\frac{(1+\sigma)}{2}\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
Moreover, $\bar{\xi}\rightarrow\phi_{\epsilon,\bar{\xi}}$ is a $C^{1}$ map.
For $\bar{\xi}\in M^{K}$, recall that $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})$, $W_{\epsilon,\xi_{i}}$ and $V_{\epsilon,\xi_{i}}$ being defined in (7) and (21), respectively. We define the functional $\bar{J}_{\epsilon}:D_{\epsilon,\rho}^{K}\subset M^{K}\rightarrow{\mathbb{R}}$, by
(46) $\bar{J}_{\epsilon}(\bar{\xi}):=J_{\epsilon}\left(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}\right)$
where $\phi_{\epsilon,\bar{\xi}}$ is given by Proposition 3.1.
The following gives the expansion of the energy functional. Its proof is presented in section 4.
Proposition 3.2.
It holds
(47) $\bar{J}_{\epsilon}(\bar{\xi})=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})+\epsilon^{4}\sum_{i=1}^{K}\Phi(\xi_{i})-\frac{1}{2}\sum_{i,j=1,i\neq j}^{K}\gamma_{ij}U\left(\frac{exp_{\xi_{i}}^{-1}\xi_{j}}{\epsilon}\right)+o(\epsilon^{4})$
$C^{0}$ uniformly with respect to $\bar{\xi}$ in compact sets of $D_{\epsilon,\rho}^{K}$ as $\epsilon$ goes to zero.
Here,
(48) $\alpha:=\frac{1}{2}\int_{{\mathbb{R}}^{n}}|\nabla U(z)|^{2}dz+\frac{1}{2}\int_{{\mathbb{R}}^{n}}U^{2}(z)dz-\frac{1}{p}\int_{{\mathbb{R}}^{n}}U^{p}(z)dz$
(49) $\beta:=\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}(z)dz-\frac{1}{n(n+2)}\int_{{\mathbb{R}}^{n}}|\nabla U(z)|^{2}|z|^{2}dz$
(50) $\gamma_{ij}:=\int_{{\mathbb{R}}^{n}}U^{p-1}(z)e^{\langle b_{ij},z\rangle}dz$
(51) $b_{ij}:=\lim_{\epsilon\rightarrow 0}\frac{exp_{\xi_{i}}^{-1}\xi_{j}}{|exp_{\xi_{i}}^{-1}\xi_{j}|}$
[…]
where $c_{1},c_{6},c_{7},c_{8}$ and $c_{9}$ are constants that depend on the dimensions $m,n$, and are given in equations (69) through (74).
Note that, in particular, from (51): $|b_{ij}|=1$. Note next that critical points of $\bar{J}_{\epsilon}$ are solutions to problem (44), by the following.
Proposition 3.3.
If $\bar{\xi}$ is an isolated, stable, $C^{1}$ critical point of $\bar{J}_{\epsilon}$, then the function $Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}$ is a solution to equation (44), or, equivalently, to problem (3).
We present the proof of this proposition in section 6. We now prove Theorem 1.1.
Proof.
(of Theorem 1.1)
Let $\bar{\xi}_{\epsilon}\in\overline{D^{K}_{\epsilon,\rho}}$ be a solution to the maximizing problem
(53) $\bar{J}_{\epsilon}(\bar{\xi}_{\epsilon})=\max\{\bar{J}_{\epsilon}(\bar{\xi}):\bar{\xi}\in\overline{D^{K}_{\epsilon,\rho}}\}$
We claim that, in fact, $\bar{\xi}_{\epsilon}\in D^{K}_{\epsilon,\rho}$.
Note that since $\xi_{0}$ of the hypothesis is a stable, critical, isolated point of $\Phi$, we may assume that is a local maximum of $\Phi$. If it were a minimum, the proof would be the same, nevertheless, we would look for a solution to a minimizing problem in eq. (53).
We first note that by Proposition 3.2, eq. (47), for $\bar{\xi}_{\epsilon}$:
(54) $\bar{J}_{\epsilon}(\bar{\xi}_{\epsilon})=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}({\xi_{\epsilon}}_{i})+\epsilon^{4}\sum_{i=1}^{K}\Phi({\xi_{\epsilon}}_{i})-\frac{1}{2}\sum_{i,j=1,i\neq j}^{K}\gamma_{ij}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)+o(\epsilon^{4})$
We now construct a particular $\bar{\eta}_{\epsilon}\in D^{K}_{\epsilon,\rho}$, so that we can make some estimates. Let $\bar{\eta}_{\epsilon}=(\eta_{1},\eta_{2},\dots,\eta_{K})$, with $\eta_{i}=\eta_{i}(\epsilon)=\exp_{\xi_{0}}(\sqrt{\epsilon}e_{i})$, $i\in\{1,2\dots,K\}$, and $e_{1},e_{2},\dots,e_{K}\in{\mathbb{R}}^{n}$, $e_{i}\neq e_{j}$, for $i\neq j$.
By direct computation, $\bar{\eta}_{\epsilon}$ satisfies
  1. i)
    $d_{g}(\xi_{0},\eta_{i})=\sqrt{\epsilon}|e_{i}|$.
  2. ii)
    $d_{g}(\eta_{i},\eta_{j})=|\exp_{\eta_{i}}^{-1}\eta_{j}|=\sqrt{\epsilon}(|e_{i}-e_{j}|+o(1)).$
Moreover,
(55) $U\left(\frac{\exp_{\eta_{i}}^{-1}\eta_{j}}{\epsilon}\right)=U\left(\frac{d_{g}(\eta_{i},\eta_{j})}{\epsilon}\right)=U\left(\frac{\sqrt{\epsilon}(|e_{i}-e_{j}|+o(1))}{\epsilon}\right)=o(\epsilon^{4})$
for $\epsilon$ small enough, since $U$ is radially symetric and of exponential decay. This implies $\bar{\eta}_{\epsilon}\in D^{K}_{\epsilon,\rho}$ holds, if $\epsilon$ is small enough. By (55) and (47)
(56) $\bar{J}_{\epsilon}(\bar{\eta}_{\epsilon})=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\eta_{i})+\epsilon^{4}\sum_{i=1}^{K}\Phi(\eta_{i})+o(\epsilon^{4})$
Note that, since it $\xi_{0}$ is a critical point,
(57) $\Phi(\eta_{i})=\Phi(\xi_{0})+\frac{1}{2}\Phi^{\prime\prime}(\xi_{0})d(\xi_{0},\eta_{i})^{2}+o(\epsilon)=\Phi(\xi_{0})+\epsilon\Phi^{\prime\prime}(\xi_{0})|e_{i}|^{2}+o(\epsilon)$
$\Phi(\xi_{0})+o(1)$
Taking into account (57) and that either $\beta=0$ or $s_{g}$ is constant, (56) turns into,
(58) $\bar{J}_{\epsilon}(\bar{\eta}_{\epsilon})=K\alpha+\epsilon^{2}\frac{1}{2}\beta Ks_{g}(\xi_{0})+\epsilon^{4}K\Phi(\xi_{0})+o(\epsilon^{4})$
Now, since $\bar{\xi}_{\epsilon}$ is a maximum of $\bar{J}_{\epsilon}$ in $\overline{D^{K}_{\epsilon,\rho}}$,
(59) $\bar{J}_{\epsilon}(\bar{\xi}_{\epsilon})\geq\bar{J}_{\epsilon}(\bar{\eta}_{\epsilon})$
Using (54) on the left hand side and (58) on the right hand side, together with the assumption that $\beta=0$ or $s_{g}$ is constant, we get
(60) $K\alpha+\epsilon^{2}\frac{1}{2}\beta Ks_{g}(\xi_{0})-\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}\gamma_{ij}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)+\epsilon^{4}\sum_{i=1}^{K}\Phi({\xi_{\epsilon}}_{i})+o(\epsilon^{4})$
$\geq K\alpha+\epsilon^{2}\frac{1}{2}\beta Ks_{g}(\xi_{0})+\epsilon^{4}K\Phi(\xi_{0})+o(\epsilon^{4})$
this is
(61) $\frac{1}{2}\sum_{i,j=1,i\neq j}^{K}\gamma_{ij}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)+\epsilon^{4}\left(K\Phi(\xi_{0})-\sum_{i=1}^{K}\Phi({\xi_{\epsilon}}_{i})\right)\leq o(\epsilon^{4})$
We now fix $\rho$ small enough so that $\xi_{0}$ is the only maximum of $\Phi$ in $B_{g}(\xi_{0},\rho)$. Hence, each term of (61) is non-negative, and therefore, bounded from above by $o(\epsilon^{4})$.
This is
(62) $0\leq\epsilon^{4}\left(K\Phi(\xi_{0})-\sum_{i=1}^{K}\Phi({\xi_{\epsilon}}_{i})\right)\leq o(\epsilon^{4})$
As a consequence $0\leq K\Phi(\xi_{0})-\sum_{i=1}^{K}\Phi({\xi_{\epsilon}}_{i})=o(1)$. It follows that $\lim_{\epsilon\rightarrow 0}\Phi({\xi_{\epsilon}}_{i})=\Phi(\xi_{0})$. Hence, $\epsilon$ small enough implies
(63) $d_{g}({\xi_{\epsilon}}_{i},\xi_{0})<\rho.$
On the other hand, also by (61), and recalling that $\gamma_{ij}=\int_{{\mathbb{R}}^{n}}U^{p-1}(z)e^{\langle b_{ij},z\rangle}dz$ and $|b_{ij}|=1$, for all $i,j\leq K$
(64) $o(\epsilon^{4})\geq\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}\gamma_{ij}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)\geq\gamma\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)$
where $\gamma:=\min\left\{\int_{{\mathbb{R}}^{n}}U^{p-1}(z)e^{\langle b_{ij},z\rangle}dz|b_{ij}\in{\mathbb{R}}^{n},|b_{ij}|=1\right\}>0.$
It follows that for $\epsilon$ small enough
(65) $\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}U\left(\frac{exp_{{\xi_{\epsilon}}_{i}}^{-1}{\xi_{\epsilon}}_{j}}{\epsilon}\right)<\epsilon^{4}.$
Finally, note that inequations (63) and (65) imply the claim that $\bar{\xi}_{\epsilon}\in D^{K}_{\epsilon,\rho}$.
4. Asymptotic expansion of the energy functional
In this section we prove Proposition 3.2. We begin with the following.
Lemma 4.1.
For $\bar{\xi}\in D^{K}_{\epsilon,\rho}$, it holds
(66) $J_{\epsilon}(Y_{\epsilon,\bar{\xi}})=J_{\epsilon}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}+\epsilon^{2}\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}\right)$
$K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})+\epsilon^{4}\sum_{i=1}^{K}\Phi(\xi_{i})-\frac{1}{2}\sum_{i\neq j}^{K}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)+o(\epsilon^{4})$
with
[…]
(67) $\alpha:=\frac{1}{2}\int_{{\mathbb{R}}^{n}}|\nabla U(z)|^{2}dz+\frac{1}{2}\int_{{\mathbb{R}}^{n}}U^{2}(z)dz-\frac{1}{p}\int_{{\mathbb{R}}^{n}}U^{p}(z)dz$
(68) $\beta:=\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}(z)dz-\frac{1}{n(n+2)}\int_{{\mathbb{R}}^{n}}|\nabla U(z)|^{2}|z|^{2}dz$
(69) $c_{1}:=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(z)}{|z|}\right)^{2}z_{1}^{4}dz\textnormal{, \ \ \ }c_{2}:=\int_{{\mathbb{R}}^{n}}U^{2}(z)z_{1}^{2}dz,$
(70) $c_{3}:=\frac{1}{54}\int_{{\mathbb{R}}^{n}}\psi(z)\frac{U^{\prime}(z)}{|z|}z_{1}^{4}dz,\textnormal{\ \ \ \ \ }c_{4}:=-\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}U(z)\psi(z)z_{1}^{2}dz$
(71) $c_{5}:=\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
(72) $c_{6}:=8c_{1}-120(n+2)c_{3}\textnormal{,\ \ \ \ \ \ }c_{7}:=-c_{3}-c_{4}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$
(73) $c_{8}:=18c_{1}+30{\bf c}c_{2}(n+2)$
(74) $c_{9}:=\frac{1}{2}{\bf c}\int_{{\mathbb{R}}^{n}}U(z)\left(\frac{1}{2}U^{\prime}(z)|z|-\frac{1}{2-p}U(z)\right)dz$
where $\psi$ is the radial function discussed in Section 2, eq. (28).
Proof.
Let $\overline{W}_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}$ and $\overline{V}_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}$. By direct computation,
$J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}}+\epsilon^{2}\overline{V}_{\epsilon,\bar{\xi}})=J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}})+\epsilon^{2}J_{\epsilon}^{\prime}(\overline{W}_{\epsilon,\bar{\xi}})[\overline{V}_{\epsilon,\bar{\xi}}]+\frac{1}{2}\epsilon^{4}J_{\epsilon}^{\prime\prime}(\overline{W}_{\epsilon,\bar{\xi}})[\overline{V}_{\epsilon,\bar{\xi}},\overline{V}_{\epsilon,\bar{\xi}}]+o(\epsilon^{5})$
(75) $J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}})+\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}\overline{W}_{\epsilon,\bar{\xi}}+(1+\epsilon^{2}\mathbf{c}s_{g})\overline{W}_{\epsilon,\bar{\xi}}-|\overline{W}_{\epsilon,\bar{\xi}}|^{p-1}\right)\overline{V}_{\epsilon,\bar{\xi}}\ d\mu_{g}$
$\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}\overline{V}_{\epsilon,\bar{\xi}}+(1+\epsilon^{2}\mathbf{c}s_{g})\overline{V}_{\epsilon,\bar{\xi}}-(p-1)|\overline{W}_{\epsilon,\bar{\xi}}|^{p-2}\overline{V}_{\epsilon,\bar{\xi}}\right)\overline{V}_{\epsilon,\bar{\xi}}\ d\mu_{g}+o(\epsilon^{5})$
Let $\mathcal{F}$ denote the last terms of (75), i.e., $J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}}+\epsilon^{2}\overline{V}_{\epsilon,\bar{\xi}})=J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}})+\mathcal{F}$.
We first compute $J_{\epsilon}\left(\overline{W}_{\epsilon,\bar{\xi}}\right)$:
$J_{\epsilon}(\overline{W}_{\epsilon,\bar{\xi}})=J_{\epsilon}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right)$
$\frac{1}{\epsilon^{n}}\int_{M}\left[\frac{1}{2}\epsilon^{2}|\nabla_{g}\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{2}+\frac{1}{2}(\epsilon^{2}\mathbf{c}s_{g}+1)\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right)^{2}-\frac{1}{p}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right)^{p}\right]d\mu_{g}$
$\frac{1}{\epsilon^{n}}\sum_{i=1}^{K}\left[\int_{M}\frac{1}{2}\epsilon^{2}|\nabla_{g}W_{\epsilon,\xi_{i}}|^{2}d\mu_{g}+\int_{M}\frac{1}{2}(\epsilon^{2}\mathbf{c}s_{g}+1)\left(W_{\epsilon,\xi_{i}}\right)^{2}d\mu_{g}-\int_{M}\frac{1}{p}\left(W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}\right]$
$\frac{1}{\epsilon^{n}}\sum_{i<j}^{K}\left[\int_{M}\epsilon^{2}\nabla_{g}W_{\epsilon,\xi_{i}}\nabla_{g}W_{\epsilon,\xi_{j}}d\mu_{g}+\int_{M}(\epsilon^{2}\mathbf{c}s_{g}+1)W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}-\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$\frac{1}{\epsilon^{n}}\left[\frac{1}{p}\int_{M}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}-\frac{1}{p}\sum_{i=1}^{K}\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}-\sum_{i<j}^{K}\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$I_{1}+I_{2}+I_{3}$
We now compute $I_{1}$:
$I_{1}=\frac{1}{\epsilon^{n}}\sum_{i=1}^{K}\left[\int_{M}\frac{1}{2}\epsilon^{2}|\nabla_{g}W_{\epsilon,\xi_{i}}|^{2}d\mu_{g}+\int_{M}\frac{1}{2}\left(W_{\epsilon,\xi_{i}}\right)^{2}d\mu_{g}-\int_{M}\frac{1}{p}\left(W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}\right]$
$\frac{1}{\epsilon^{n}}\sum_{i=1}^{K}\int_{M}\frac{1}{2}(\epsilon^{2}\mathbf{c}s_{g})\left(W_{\epsilon,\xi_{i}}\right)^{2}d\mu_{g}$
$A_{1}+A_{2}$
It was computed in [8], eq. (56), that
(76) $A_{1}=K\alpha+\epsilon^{2}\sum_{i=1}^{K}\Theta_{1}(\xi_{i})+\epsilon^{4}\sum_{i=1}^{K}\Theta_{2}(\xi_{i})+o(\epsilon^{4})$
with
(77) $\Theta_{1}(\xi)=-c_{1}s_{g}(\xi),$
and
[…]
where $Ric_{\xi}$ is the Ricci curvature and $R_{\xi}$ the curvature tensor, meanwhile $\alpha$ and $c_{1}$ are dimensional constants that depend only on $n$ and $m$, given by eqs. (67) and (69).
We now estimate
$A_{2}=\epsilon^{2}\mathbf{c}\frac{1}{2}\sum_{i=1}^{K}\frac{1}{\epsilon^{n}}\int_{M}s_{g}W_{\epsilon,\xi_{i}}^{2}d\mu_{g}.$
Consider Taylor’s expansions of $s_{g}$ around $\xi$,
(79) $s_{g}(\exp_{\xi}(\epsilon z))=s_{g}(\xi)+\epsilon\partial_{j}s_{g}(\xi)z_{j}+\epsilon^{2}\frac{1}{2}\partial^{2}_{lj}s_{g}(\xi)z_{l}z_{j}+o(\epsilon^{2})$
and recall the expansions of $g$ and $det(g)$ around $\xi$ (Lemma 2.1). We get
$\frac{1}{\epsilon^{n}}\frac{1}{2}\int_{M}s_{g}W_{\epsilon,\xi_{i}}^{2}d\mu_{g}=\frac{1}{2}\int_{B(0,r/\epsilon)}s_{g}(\exp_{\xi}(\epsilon z))(U(z)\chi_{r/\epsilon}(\epsilon z))^{2}\sqrt{det(g_{\xi}(\epsilon z))}dz$
$\frac{1}{2}s_{g}(\xi_{i})\int_{{\mathbb{R}}^{n}}U^{2}dz-\frac{\epsilon^{2}}{12}s_{g}(\xi_{i})R_{kl}\int_{{\mathbb{R}}^{n}}U^{2}z_{k}z_{l}dz$
$\epsilon\frac{1}{2}\partial_{j}s_{g}(\xi_{i})\int_{{\mathbb{R}}^{n}}U^{2}z_{j}dz+\epsilon^{2}\frac{1}{4}\partial^{2}_{lj}s_{g}(\xi_{i})\int_{{\mathbb{R}}^{n}}U^{2}z_{j}z_{l}dz+o(\epsilon^{2})$
Now, since $U^{2}$ is radial, for $i,j\in\{1,2,\dots,n\}$,
$\int_{{\mathbb{R}}^{n}}U^{2}z_{j}z_{l}dz=\delta_{jl}\int_{{\mathbb{R}}^{n}}U^{2}z_{j}z_{l}dz=\delta_{jl}\int_{{\mathbb{R}}^{n}}U^{2}z_{1}^{2}dz$
Also, for any $j\in\{1,2,\dots,n\}$, $\int_{{\mathbb{R}}^{n}}U^{2}z_{j}dz=0$, since $U$ is radial. This implies
(80) $A_{2}=\epsilon^{2}\frac{1}{2}\mathbf{c}\left(\sum_{i=1}^{K}s_{g}(\xi_{i})\right)\int_{{\mathbb{R}}^{n}}U^{2}dz-\frac{\epsilon^{4}}{4}\mathbf{c}\left(\frac{1}{3}\sum_{i=1}^{K}s_{g}(\xi_{i})^{2}+\sum_{i=1}^{K}\Delta s_{g}(\xi_{i})\right)\int_{{\mathbb{R}}^{n}}U^{2}z_{1}^{2}dz+o(\epsilon^{4})$
Note that adding the terms of order $\epsilon^{2}$ from $A_{1}$ and $A_{2}$ gives
$\frac{1}{2}\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}dz\sum_{i=1}^{K}s_{g}(\xi_{i})-c_{1}\sum_{i=1}^{K}s_{g}(\xi_{i})$
$\left(\frac{1}{2}\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}dz-\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz\right)\sum_{i=1}^{K}s_{g}(\xi_{i})$
$\frac{1}{2}\left(\mathbf{c}\int_{{\mathbb{R}}^{n}}U^{2}(z)-\frac{1}{n(n+2)}\int|\nabla U(z)|^{2}|z|^{2}dz\right)\sum_{i=1}^{K}s_{g}(\xi_{i})=\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})$
where we have used the equality (from [19], page 14):
$\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz=\frac{3}{n(n+2)}\int_{{\mathbb{R}}^{n}}|\nabla U|^{2}|z|^{2}dz$
and $\beta$ is the dimensional constant from equation (68). We conclude
(81) $I_{1}=A_{1}+A_{2}=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})+\epsilon^{4}\left(\sum_{i=1}^{K}\Theta_{2}(\xi_{i})-\frac{\mathbf{c}}{4}c_{2}\left(\frac{1}{3}\sum_{i=1}^{K}s_{g}(\xi_{i})^{2}+\sum_{i=1}^{K}\Delta s_{g}(\xi_{i})\right)\right)+o(\epsilon^{4})$
with $c_{2}:=\int_{{\mathbb{R}}^{n}}U^{2}z_{1}^{2}dz$. We now estimate $I_{2}$,
$I_{2}=\frac{1}{\epsilon^{n}}\sum_{i<j}^{K}\left[\int_{M}\epsilon^{2}\nabla_{g}W_{\epsilon,\xi_{i}}\nabla_{g}W_{\epsilon,\xi_{j}}d\mu_{g}+\int_{M}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}-\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$\epsilon^{2}\mathbf{c}\sum_{i<j}^{K}\frac{1}{\epsilon^{n}}\int_{M}s_{g}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}=:B_{1}+B_{2}$
It is shown in [4], eq. (4.4) that for $i\neq j$,
$B_{1}=\frac{1}{\epsilon^{n}}\sum_{i<j}^{K}\left[\int_{M}\epsilon^{2}\nabla_{g}W_{\epsilon,\xi_{i}}\nabla_{g}W_{\epsilon,\xi_{j}}d\mu_{g}+\int_{M}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}-\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$o(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})$
which is of order $o(\epsilon^{4})$, for $\xi_{i},\xi_{j}$ components of $\bar{\xi}\in D^{K}_{\epsilon,\rho}$, by (38).
On the other hand, since $s_{g}$ is bounded, for some $C_{s}$ upper bound for $s_{g}$:
$\epsilon^{2}\mathbf{c}\sum_{i<j}^{K}\frac{1}{\epsilon^{n}}\int_{M}s_{g}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}<\epsilon^{2}C_{s}\mathbf{c}\sum_{i<j}^{K}\frac{1}{\epsilon^{n}}\int_{M}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}$
Let $\exp_{\xi_{i}}^{-1}(x)=\epsilon z$. Then, for some $C>0$, $\delta>0$ small, $i\neq j$,
$\frac{1}{\epsilon^{n}}\int_{M}W_{\epsilon,\xi_{i}}W_{\epsilon,\xi_{j}}d\mu_{g}$
$\int_{B(0,r/\epsilon)}U(z)\chi_{r}(z)U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)\chi_{r}(\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z)))|g_{\xi_{i}}(\epsilon z)|dz$
$\leq C\int_{B(0,r/\epsilon)}U(z)U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)dz=o(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})$
where in the last equality we have used an estimation from Lemma 6.2. Note that $o\left(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)=o(\epsilon^{4})$, for $\xi_{i},\xi_{j}$ components of $\bar{\xi}\in D^{K}_{\epsilon,\rho}$, by (38). This implies $I_{2}=o(\epsilon^{4})$. We now claim
(82) $I_{3}=-\frac{1}{2}\sum_{i\neq j}^{K}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)+o(\epsilon^{4})$
where $\gamma_{ij}$ is defined in (50).
It is shown in [4], eq. (4.7) that
$I_{3}=-\frac{1}{\epsilon^{n}}\left[\frac{1}{p}\int_{M}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}-\frac{1}{p}\sum_{i=1}^{K}\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p}d\mu_{g}-\sum_{i<j}^{K}\int_{M}\left(W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$\frac{1}{\epsilon^{n}}\int_{M}\sum_{j=1}^{K-1}\left(\sum_{i=j+1}^{K}W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}+o(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})$
and in eq. (4.8), (4.10) and (4.11) of the same article, that
$\frac{1}{\epsilon^{n}}\int_{M}\left(\sum_{i=j+1}^{K}W_{\epsilon,\xi_{i}}\right)^{p-1}W_{\epsilon,\xi_{j}}d\mu_{g}=\sum_{i\neq j}^{K}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)+o\left(e^{-p\frac{d_{g}(\xi_{i},\xi_{j})}{2\epsilon}}\right)$
Taking into account that $o(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})=o(\epsilon^{4})$ and $o(e^{-p\frac{d_{g}(\xi_{i},\xi_{j})}{2\epsilon}})=o(\epsilon^{4})$, by (38), this implies the claim of eq. (82).
We now compute the terms corresponding to $\mathcal{F}$, from eq. (75).
$\mathcal{F}:=\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}-|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-1}\right)\sum_{j=1}^{K}V_{\epsilon,\xi_{j}}\ d\mu_{g}$
$\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}-(p-1)|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-2}\sum_{l=1}^{K}V_{\epsilon,\xi_{l}}\right)\sum_{j=1}^{K}V_{\epsilon,\xi_{j}}\ d\mu_{g}+o(\epsilon^{5})$
$\sum_{i=1}^{K}\left[\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}W_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})W_{\epsilon,\xi_{i}}-|W_{\epsilon,\xi_{i}}|^{p-1}\right)V_{\epsilon,\xi_{i}}\ d\mu_{g}\right]$
$\sum_{i=1}^{K}\left[\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\epsilon^{2}\Delta_{g}V_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})V_{\epsilon,\xi_{i}}-(p-1)|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}\right)V_{\epsilon,\xi_{i}}\ d\mu_{g}\right]$
$\epsilon^{2}\frac{1}{\epsilon^{n}}\sum_{i\neq j}^{K}\left[\int_{M}\epsilon^{2}\nabla W_{\epsilon,\xi_{i}},\nabla V_{\epsilon,\xi_{j}}d\mu_{g}+(1+\epsilon^{2}\mathbf{c}s_{g})W_{\epsilon,\xi_{i}}V_{\epsilon,\xi_{j}}d\mu_{g}-|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\sum_{i\neq j}^{K}\int_{M}\epsilon^{2}\nabla V_{\epsilon,\xi_{i}},\nabla V_{\epsilon,\xi_{j}}d\mu_{g}+\int_{M}(1+\epsilon^{2}\mathbf{c}s_{g})V_{\epsilon,\xi_{i}}V_{\epsilon,\xi_{j}}d\mu_{g}-\int_{M}(p-1)|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}V_{\epsilon,\xi_{j}}d\mu_{g}$
$\epsilon^{2}\frac{1}{\epsilon^{n}}\left[\sum_{j=1}^{K}\int_{M}-|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}+\sum_{i=1}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{i}}\ d\mu_{g}+\sum_{i\neq j}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$\epsilon^{4}\frac{(p-1)}{2\epsilon^{n}}\left[-\sum_{j,l=1}^{K}\int_{M}|\sum_{i\neq j}^{K}W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}+\sum_{i=1}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}+\sum_{i\neq j}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}V_{\epsilon,\xi_{j}}d\mu_{g}\right]$
$o(\epsilon^{5})$
$K_{1}+K_{2}+K_{3}+K_{4}+K_{5}+K_{6}+o(\epsilon^{5})$
We first consider $K_{1}+K_{2}$. Note that
$K_{1}+K_{2}=\sum_{i=1}^{K}\left[\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi_{i}}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi_{i}}-|W_{\epsilon,\xi_{i}}|^{p-1}\right)V_{\epsilon,\xi_{i}}\ d\mu_{g}\right]$
$\sum_{i=1}^{K}\left[\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}V_{\epsilon,\xi_{i}}+V_{\epsilon,\xi_{i}}-(p-1)|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}\right)V_{\epsilon,\xi_{i}}\ d\mu_{g}\right]$
$\sum_{i=1}^{K}\ \epsilon^{6}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\mathbf{c}s_{g}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}+o(\epsilon^{4})$
and since $V_{\epsilon,\xi_{i}}$ is bounded and exponentially decreasing,
$\epsilon^{6}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\mathbf{c}s_{g}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}=o(\epsilon^{5})$
Hence
(83) $K_{1}+K_{2}=\sum_{i=1}^{K}\mathcal{A}(\xi_{i})+o(\epsilon^{4})$
where
$\mathcal{A}(\xi):=\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi}-|W_{\epsilon,\xi}|^{p-1}\right)V_{\epsilon,\xi}\ d\mu_{g}$
(84) $\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}V_{\epsilon,\xi}+V_{\epsilon,\xi}-(p-1)|W_{\epsilon,\xi}|^{p-2}V_{\epsilon,\xi}\right)V_{\epsilon,\xi}d\mu_{g}+o(\epsilon^{5})$
Note that by construction of $V$, eq. (21), and using eq. (6) for $U$, the term
$\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi}-|W_{\epsilon,\xi}|^{p-1}\right)V_{\epsilon,\xi}\ d\mu_{g}$
is of order $\epsilon^{2}$, more specifically
(85) $\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi}-|W_{\epsilon,\xi}|^{p-1}\right)V_{\epsilon,\xi}\ d\mu_{g}=-\epsilon^{4}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
Meanwhile, using the fact that $\Delta_{g}V-\Delta_{{\mathbb{R}}^{n}}V=o(\epsilon^{2})$ and rewriting the second term of (84) on ${\mathbb{R}}^{n}$:
(86) $\epsilon^{4}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}V_{\epsilon,\xi}+V_{\epsilon,\xi}-(p-1)|W_{\epsilon,\xi}|^{p-2}V_{\epsilon,\xi}\right)V_{\epsilon,\xi}d\mu_{g}=\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
Hence, by adding (85) and (86), (84) turns into
(87) $\mathcal{A}(\xi)=-\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
Recall also that $V=V_{1}+V_{2}$, with $V_{1}$ and $V_{2}$ given by eqs. (29) and (31). Hence
(88) $\int_{{\mathbb{R}}^{n}}L_{0}VVdz=\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz+\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{2}dz+\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{1}dz+\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{2}dz$
The first term, $\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz$, can be computed as follows, using eq. (27) for $L_{0}V_{1}$ and eq. (29) for $V_{1}$,
$\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz=-\epsilon^{4}\frac{1}{2}\frac{1}{9}\left(R_{kl}R_{ij}\int_{{\mathbb{R}}^{n}}\psi\frac{U^{\prime}}{|z|}z_{k}z_{l}z_{i}z_{j}dz\right)$
[…]
since $\psi\frac{U^{\prime}}{|z|}$ is radial; with $c_{3}$ a dimensional constant, given by
(90) $c_{3}:=\frac{1}{54}\int_{{\mathbb{R}}^{n}}\psi(|z|)\frac{U^{\prime}(|z|)}{|z|}z_{1}^{4}dz,$
where $\psi=\psi(|z|)$ is the radial function discussed in Section 2, eqs. (28) and (29). Note that for (89), we have used the fact that for any radial function $\varphi(|z|)$ in ${\mathbb{R}}^{n}$,
$\int_{\mathbb{R}^{n}}\varphi(|z|)z_{i}z_{j}z_{k}z_{l}dz=(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{R}^{n}}\varphi(|z|)z_{1}^{2}z_{2}^{2}$
$\frac{1}{3}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{R}^{n}}\varphi(|z|)z_{1}^{4}$
which follows from the fact that
$\int_{\mathbb{S}^{n-1}}z_{i}z_{j}z_{k}z_{l}=(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{S}^{n-1}}z_{1}^{2}z_{2}^{2}$
$\frac{1}{3}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{S}^{n-1}}z_{1}^{4}$
In a similar way, for any radial function $\varphi(|z|)$ on ${\mathbb{R}}^{n}$,
(91) $\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{k}z_{l}dz=\delta_{kl}\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{1}^{2}dz$
We now compute the rest of the terms in (88). For the second term in (88), by direct computation, using eq. (27) for $L_{0}V_{1}$ and eq. (31) for $V_{2}$,
$\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{2}dz=\frac{1}{3}R_{kl}{\bf c}s_{g}\int_{{\mathbb{R}}^{n}}\frac{U^{\prime}(z)}{|z|}\left(\frac{1}{2}U^{\prime}(z)|z|-\frac{1}{2-p}U(z)\right)z_{k}z_{l}dz$
$\frac{{\bf c}}{3}s_{g}^{2}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
by (91), since $\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)$ is a radial function. On the other hand, by eq. (26) for $L_{0}V_{2}$ and eq. (29) for $V_{1}$,
$\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{1}dz=-\frac{1}{3}{\bf c}s_{g}R_{kl}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{k}z_{l}dz$
$\frac{1}{3}{\bf c}s_{g}R_{kl}\delta_{kl}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz=-\frac{1}{3}{\bf c}s_{g}^{2}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz=2c_{4}s_{g}^{2}$
with
$c_{4}=-\frac{1}{6}{\bf c}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz$
since $U(z)\psi(|z|)$ a radial function. And by eq. (26) for $L_{0}V_{2}$ and eq. (31) for $V_{2}$:
$\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{2}dz=-{\bf c}s_{g}\int_{{\mathbb{R}}^{n}}U(z)\left(\frac{1}{2}U^{\prime}(z)|z|-\frac{1}{2-p}U(z)\right)dz$
We then get, from these equations and (83), together with (87) and (88),
$K_{1}+K_{2}=\mathcal{A}(\xi)+o(\epsilon^{4})=-\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
[…]
with dimensional constants $c_{5}$, $c_{9}$ given by,
(93) $c_{5}:=\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
(94) $c_{9}:=\frac{1}{2}{\bf c}\int_{{\mathbb{R}}^{n}}U(z)\left(\frac{1}{2}U^{\prime}(z)|z|-\frac{1}{2-p}U(z)\right)dz$
On the other hand, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ are $o(\epsilon^{4})$ by Lemmas 6.3, 6.4, 6.6 and 6.7, respectively. We conclude,
$J_{\epsilon}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}+\epsilon^{2}\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}\right)=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})+\epsilon^{4}\sum_{i=1}^{K}\Phi(\xi_{i})-\frac{1}{2}\sum_{i\neq j}^{K}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)+o(\epsilon^{4})$
with
[…]
where, $c_{6}=8c_{1}-120(n+2)c_{3}$, $c_{7}=-c_{3}-c_{4}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$ and $c_{8}=18c_{1}+30{\bf c}c_{2}(n+2)$.
We now are ready to prove Proposition 3.2.
Proof.
(of Proposition 3.2)
Recall that $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})$. We will prove that
(95) $J_{\epsilon}\left(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}\right)-J_{\epsilon}\left(Y_{\epsilon,\bar{\xi}}\right)=o(\epsilon^{4})$
and the Proposition will follow from Lemma 4.1. We write $F(u)=\frac{1}{p}(u^{+})^{p}$ and $f(u)=(u^{+})^{p-1}$. First, since $\phi_{\epsilon,\bar{\xi}}\in K^{\perp}_{\epsilon,\bar{\xi}}$,
$0=\langle\phi_{\epsilon,\bar{\xi}},S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})\rangle_{\epsilon}=\langle\phi_{\epsilon,\bar{\xi}},(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})-i^{*}_{\epsilon}(f(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}))\rangle_{\epsilon}$
[…]
Hence, by direct computation and the last equation,
[…]
$\frac{1}{\epsilon^{n}}\int_{M}[F(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})-F(Y_{\epsilon,\bar{\xi}})-f(Y_{\epsilon,\bar{\xi}})\phi_{\epsilon,\bar{\xi}}]d\mu_{g}$
[…]
Since, for some $t_{1},t_{2}\in[0,1]$, by the mean value Theorem,
$\frac{1}{\epsilon^{n}}\int_{M}[f(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})-f(Y_{\epsilon,\bar{\xi}})]\phi_{\epsilon,\bar{\xi}}d\mu_{g}=\frac{1}{\epsilon^{n}}\int_{M}f^{\prime}(Y_{\epsilon,\bar{\xi}}+t_{1}\phi_{\epsilon,\bar{\xi}})\phi^{2}_{\epsilon,\bar{\xi}}d\mu_{g}$
and
$\frac{1}{\epsilon^{n}}\int_{M}[F(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})-F(Y_{\epsilon,\bar{\xi}})-f(Y_{\epsilon,\bar{\xi}})\phi_{\epsilon,\bar{\xi}}]d\mu_{g}=\frac{1}{2\epsilon^{n}}\int_{M}f^{\prime}(Y_{\epsilon,\bar{\xi}}+t_{2}\phi_{\epsilon,\bar{\xi}})\phi^{2}_{\epsilon,\bar{\xi}}d\mu_{g}$
Moreover, for any $t\in[0,1]$,
$\frac{1}{\epsilon^{n}}\int_{M}f^{\prime}(Y_{\epsilon,\bar{\xi}}+t\phi_{\epsilon,\bar{\xi}})\phi^{2}_{\epsilon,\bar{\xi}}d\mu_{g}\leq c\frac{1}{\epsilon^{n}}\int_{M}Y_{\epsilon,\bar{\xi}}^{p-2}\phi_{\epsilon,\bar{\xi}}^{2}d\mu_{g}+c\frac{1}{\epsilon^{n}}\int_{M}\phi_{\epsilon,\bar{\xi}}^{p}d\mu_{g}$
[…]
where, we have used, from eq. (45), that $\phi_{\epsilon,\bar{\xi}}||_{\epsilon}=o(\epsilon^{2})$, since $\bar{\xi}\in D^{K}_{\epsilon,\rho}$.
5. The Finite dimensional reduction
In this section we sketch a proof for the finite dimensional reduction, Prop. 3.1. We follow [4, 7, 19], where more detailed proofs in similar situations can be found.
[…]
$0=\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi)\right\}=\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}})+S^{\prime}_{\epsilon}(Y_{\epsilon,\bar{\xi}})\phi+\bar{N}_{\epsilon,\bar{\xi}}(\phi)\right\}=-R_{\epsilon,\bar{\xi}}+L_{\epsilon,\bar{\xi}}-N_{\epsilon,\bar{\xi}}(\phi)$
with
(96) $R_{\epsilon,\bar{\xi}}:=-\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}})\right\}=\Pi_{\epsilon,\xi}^{\perp}\left\{i^{*}_{\epsilon}(f(Y_{\epsilon,\bar{\xi}}))-Y_{\epsilon,\bar{\xi}}\right\}$
(97) $L_{\epsilon,\bar{\xi}}(\phi):=\Pi_{\epsilon,\xi}^{\perp}\left\{S^{\prime}_{\epsilon}(Y_{\epsilon,\bar{\xi}})\phi\right\}=\Pi_{\epsilon,\xi}^{\perp}\left\{\phi-i^{*}_{\epsilon}(f^{\prime}(Y_{\epsilon,\bar{\xi}})\phi)\right\}$
(98) $N_{\epsilon,\bar{\xi}}(\phi):=\Pi_{\epsilon,\xi}^{\perp}\left\{\bar{N}_{\epsilon,\bar{\xi}}(\phi)\right\}=\Pi_{\epsilon,\xi}^{\perp}\left\{i^{*}_{\epsilon}\left(f(Y_{\epsilon,\bar{\xi}}+\phi)-f(Y_{\epsilon,\bar{\xi}})-f^{\prime}(Y_{\epsilon,\bar{\xi}})\phi\right)\right\}$
In this way, we can rewrite eq. (43) as $L_{\epsilon,\bar{\xi}}(\phi)=N_{\epsilon,\bar{\xi}}(\phi)+R_{\epsilon,\bar{\xi}}$.
Moreover, if $L_{\epsilon,\bar{\xi}}$ is invertible, the problem in eq. (43) turns into a fixed point problem for the operator,
$T_{\epsilon,\bar{\xi}}(\phi)=L^{-1}_{\epsilon,\bar{\xi}}\left(N_{\epsilon,\bar{\xi}}(\phi)+R_{\epsilon,\bar{\xi}}\right)$
Following this idea, we begin by proving that $L_{\epsilon,\bar{\xi}}$ is invertible, for suitable $\bar{\xi}$ and $\epsilon$.
Lemma 5.1.
There is some $\epsilon_{0}$, $c>0$, such that for any $\epsilon\in(0,\epsilon_{0})$ and $\bar{\xi}\in M^{K}$, $\bar{\xi}=(\xi_{1},\xi_{2},\dots,\xi_{K})$, such that if
(99) $\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}U\left(\frac{exp_{\xi_{i}}^{-1}\xi_{j}}{\epsilon}\right)<\epsilon^{4},$
[…]
Proof.
We follow [12, Proposition 3.1], [19, Lemma 6.1]. We proceed by contradiction. Suppose there is a sequence $\epsilon_{j}\rightarrow 0$, $\{\epsilon_{j}\}_{j\in{\mathbb{N}}}$ together with $\{\bar{\xi}_{j}\}_{j\in{\mathbb{N}}}\subset M^{K}$, such that the hypothesis of the Lemma, eq. (99) is satisfied and at the same time there is some $\{\phi_{j}\}_{j\in{\mathbb{N}}}\subset K^{\perp}_{\epsilon,\bar{\xi}}$, with
[…]
where $\psi_{j}:=L_{\epsilon,\bar{\xi}}(\phi_{j})$. Hence
(101) $\phi_{j}-i^{*}_{\epsilon_{j}}(f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})\phi_{j})=\psi_{j}+\zeta_{j}$
with $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}\left(U_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)$ and $\zeta_{j}:=\Pi_{\epsilon,\xi}\left\{\phi_{j}-i^{*}_{\epsilon_{j}}(f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})\phi_{j})\right\}$. Let $w_{j}=\phi_{j}-(\psi_{j}+\zeta_{j})$. We will prove that the existence of such sequences lead to the following contradiction
(102) $\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}\rightarrow 1$
and
(103) $\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}\rightarrow 0$
We first prove eq. (102). Being $\zeta_{j}\in K_{\epsilon_{j},\bar{\xi}_{j}}$ we write $\zeta_{j}=\sum_{i=1}^{K}\sum_{m=1}^{n}a_{j}^{mi}Z_{\epsilon_{j},\bar{\xi}_{j}}^{m}$. For $m\in\{1,2,\dots,n\}$ and $l\in\{1,2,\dots,K\}$, we multiply $\psi_{j}+\zeta_{j}$ by $Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}$, $h\in\{1,2,\dots,n\}$ (using eq. (101))
(104) $\sum_{i=1}^{K}\sum_{m=1}^{n}a_{j}^{mi}\langle Z_{\epsilon_{j},\bar{\xi}_{j}}^{m},Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\rangle_{\epsilon_{j}}=\langle\phi_{j},Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\rangle_{\epsilon_{j}}-\langle i^{*}_{\epsilon_{j}}(f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})\phi_{j}),Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\rangle_{\epsilon_{j}}$
on the other hand, by (149)
(105) $\sum_{i=1}^{K}\sum_{m=1}^{n}a_{j}^{mi}\langle Z_{\epsilon_{j},\bar{\xi}_{j}}^{m},Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\rangle_{\epsilon_{j}}=Ca_{j}^{hl}+o(1)$
Hence, by (104) and (105)
(106) $Ca_{j}^{hl}+o(1)=\frac{1}{\epsilon_{j}^{n}}\int_{M}\left[\epsilon_{j}^{2}\nabla_{g}Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\nabla_{g}\phi_{j}+(\epsilon_{j}^{2}\mathbf{c}s_{g}+1)Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\phi_{j}-f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\phi_{j}\right]d\mu_{g}$
Let $\tilde{\phi_{l}}_{j}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ be given by
(107) $\tilde{\phi_{l}}_{j}(z)=\begin{cases}\phi_{lj}(exp_{{\xi_{l}}_{j}}(\epsilon_{j}z)\chi_{r}(\epsilon_{j}z))&\textit{if}\ \ z\in B(0,r/\epsilon_{j})\\ 0&\textit{otherwise}\end{cases}$
[…]
$\left|\frac{1}{\epsilon^{n}}\int_{M}\epsilon^{2}_{j}\mathbf{c}s_{g}Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}}\phi_{j}d\mu_{g}\right|\leq\epsilon_{j}^{2}c_{1}\left|\int_{B(0,r/\epsilon_{j})}\psi^{h}(z)\chi_{r}(\epsilon_{j}z)\tilde{\phi_{l}}_{j}(z)|g_{{\xi_{l}}_{j}}(\epsilon_{j}z)|^{1/2}dz\right|$
$\epsilon_{j}^{2}c_{1}\left(\int_{{\mathbb{R}}^{n}}\psi^{h}\tilde{\phi_{l}}_{j}dz+o(1)\right)\leq\epsilon_{j}^{2}c_{1}\left(\int_{{\mathbb{R}}^{n}}(\psi^{h})^{2}dz\right)^{1/2}\left(\int_{{\mathbb{R}}^{n}}\tilde{\phi_{l}}_{j}^{2}dz\right)^{1/2}+o(\epsilon_{j}^{2})$
[…]
for some constants, $c_{1}$, $c_{2}$, since $s_{g}$, $\nabla U||_{L^{2}({\mathbb{R}}^{n})}$ and $\tilde{\phi}||_{L^{2}({\mathbb{R}}^{n})}$ are bounded. Therefore
(108) $Ca_{j}^{hl}+o(1)=\frac{1}{\epsilon_{j}^{n}}\int_{M}\left[\epsilon_{j}^{2}\nabla_{g}Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\nabla_{g}\phi_{j}+(\epsilon_{j}^{2}\mathbf{c}s_{g}+1)Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\phi_{j}-f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})\phi_{j}Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\right]d\mu_{g}$
$\frac{1}{\epsilon_{j}^{n}}\int_{M}\left[\epsilon_{j}^{2}\nabla_{g}Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\nabla_{g}\phi_{j}+Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\phi_{j}-f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})\phi_{j}Z_{\epsilon_{j},\bar{\xi}_{j}}^{h}\right]d\mu_{g}+o(\epsilon_{j})$
$\int_{{\mathbb{R}}^{n}}\left(\nabla\psi^{h}\nabla\tilde{\phi}+\psi^{h}\tilde{\phi}-f^{\prime}(U+\epsilon_{j}^{2}V)\tilde{\phi}\psi^{h}dz\right)+o(1)$
$\int_{{\mathbb{R}}^{n}}\left(\nabla\psi^{h}\nabla\tilde{\phi}+\psi^{h}\tilde{\phi}-f^{\prime}(U)\psi^{h}\tilde{\phi}dz\right)+o(\epsilon_{j})+o(1)=o(1)$
It follows that $a^{hl}_{j}\rightarrow 0$ for any $h=1,2,\cdots n$ and $l=1,2,\cdots,K$. Hence
(109) $\zeta_{j}||_{\epsilon_{j}}\rightarrow 0\text{\ as \ }j\rightarrow\infty$
And then, for $w_{j}=\phi_{j}-(\psi_{j}+\zeta_{j})$, assumptions on $\psi_{j}||_{\epsilon_{j}}$, (100) and (109) imply
(110) $w_{j}||_{\epsilon_{j}}\rightarrow 1$
Moreover, $w_{j}=i^{*}(f^{\prime}(Y_{\epsilon_{j},\bar{\xi}})\phi_{j})$, from (101). Hence, $w_{j}$ satisfies weakly, on $M$,
(111) $\epsilon^{2}_{j}\Delta_{g}w_{j}+(\epsilon_{j}^{2}\mathbf{c}s_{g}+1)w_{j}=f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}+f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})(\psi_{j}+\zeta_{j})$
Multiplying eq. (111) by $w_{j}$ and integrating over $M$,
(112) $w_{j}||_{\epsilon_{j}}^{2}=\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}+\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})(\psi_{j}+\zeta_{j})w_{j}d\mu_{g}$
By Hölder’s inequality, the second term is $o(1)$:
$\left|\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})(\psi_{j}+\zeta_{j})w_{j}d\mu_{g}\right|$
[…]
since $\psi_{j}||_{\epsilon_{j}}\rightarrow 0$, $\zeta_{j}||_{\epsilon_{j}}\rightarrow 0$ and $w_{j}||_{\epsilon_{j}}\rightarrow 1$ as $j\rightarrow\infty$. From this and eq. (112), eq. (102) follows.
We now prove eq. (103). For $l\in\{1,2,\dots,K\}$, we define $\tilde{w_{l}}_{j}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$,
$\tilde{w_{l}}_{j}(z)=w_{j}(\exp_{{\xi_{l}}_{j}}(\epsilon_{j}z))\chi_{r}(\exp_{{\xi_{l}}_{j}}(\epsilon_{j}z)),z\in{\mathbb{R}}^{n}$
[…]
(113) $\Delta\tilde{w}_{l}+\tilde{w}_{l}=f^{\prime}(U)\tilde{w}_{l}$
Let $\varphi\in C_{0}({\mathbb{R}}^{n})$. Let $\varphi_{j}(x):=\varphi\left(\frac{exp^{-1}_{{\xi_{l}}_{j}}(x)}{\epsilon_{j}}\right)\chi_{r}(exp^{-1}_{{\xi_{l}}_{j}}(x))$, for $x\in B({\xi_{l}}_{j},\epsilon_{j}R)\subset M$. For some $R>0$ big enough so that $supp\ \varphi\subset B(0,R)$ and $j$ big enough such that $B({\xi_{l}}_{j},\epsilon_{j}R)\subset B({\xi_{l}}_{j},r)$.
Now, multiplying (111) by $\varphi_{j}$ and integrating over $M$
(114) $\frac{1}{\epsilon_{j}^{n}}\int_{M}(\epsilon_{j}^{2}\nabla_{g}w_{j}\nabla_{g}\varphi_{j}+(1+s_{g}\mathbf{c}\epsilon_{j}^{2})w_{j}\varphi_{j})d\mu_{g}=\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}\varphi_{j}d\mu_{g}+\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})(\psi_{j}+\zeta_{j})\varphi_{j}d\mu_{g}$
Hence, on ${\mathbb{R}}^{n}$, with $x=\exp_{{\xi_{l}}_{j}}(\epsilon_{j}z)$,
$\int_{B(0,R)}\left(\sum_{s,t=1}^{n}g_{{\xi_{l}}_{j}}^{st}(\epsilon_{j}z)\frac{\partial\tilde{w_{l}}_{j}}{\partial z_{s}}\frac{\partial\varphi}{\partial z_{t}}+(1+s_{g}\mathbf{c}\epsilon_{j}^{2})\tilde{w_{l}}_{j}\varphi\right)|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|^{1/2}dz$
$=\int_{B(0,R)}f^{\prime}\left((U(z)+\epsilon_{j}^{2}V(z))\chi_{r}(\epsilon_{j}z)+\sum_{i\neq l}Y\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z)}{\epsilon_{j}}\right)\chi_{r}(\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))\right)$
${w_{l}}_{j}\ \varphi\ |g_{{\xi_{l}}_{j}}(\epsilon_{j}z)|^{1/2}dz$
$+\int_{B(0,R)}f^{\prime}\left((U(z)+\epsilon_{j}^{2}V(z))\chi_{r}(\epsilon_{j}z)+\sum_{i\neq l}Y\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z)}{\epsilon_{j}}\right)\chi_{r}(\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))\right)$
$(\tilde{\psi}_{j}+\tilde{\zeta}_{j})\ \varphi\ |g_{{\xi_{l}}_{j}}(\epsilon_{j}z)|^{1/2}dz$
where $\tilde{\psi}_{j}:=\psi_{j}(\exp_{{\xi_{l}}_{j}}(\epsilon_{j}z))$ and $\tilde{\zeta}_{j}(z):=\zeta_{j}(\exp_{{\xi_{l}}_{j}}(\epsilon_{j}z))$ for $z\in B(0,R/\epsilon_{j})$. Also, we estimate the third term of the left-hand side to be $o(\epsilon_{j})$:
$\int_{B(0,R)}{\bf c}s_{g}\epsilon_{j}^{2}\tilde{w_{l}}_{j}\varphi|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|^{1/2}dz\leq c\epsilon_{j}^{2}\int_{B(0,R)}\tilde{w_{l}}_{j}\varphi|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|^{1/2}dz$
$\leq c\epsilon_{j}^{2}\left(\int_{B(0,R)}\tilde{w_{l}}_{j}^{2}dz\right)^{1/2}\left(\int_{B(0,R)}\varphi^{2}|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|dz\right)^{1/2}$
[…]
with $c$ an upper bound for $s_{g}$, $c_{2}$ a bound for $\int_{B(0,R)}\varphi^{2}|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|dz$ and $\tilde{w_{l}}_{j}^{2}$ being bounded in $H^{1}({\mathbb{R}}^{n})$}, independently of $j$. Hence, as $\epsilon_{j}\rightarrow 0$, from eq.(115) we get
(116) $\int_{{\mathbb{R}}^{n}}\left(\sum_{s,t=1}^{n}\delta_{st}\frac{\partial\tilde{w_{l}}_{j}}{\partial z_{s}}\frac{\partial\varphi}{\partial z_{t}}+\tilde{w_{l}}_{j}\varphi\right)dz=\int_{{\mathbb{R}}^{n}}f^{\prime}\left(U(z)\right){w_{l}}_{j}\varphi dz$
since $\tilde{\psi}_{j}$, $\tilde{\zeta}_{j}\rightarrow 0$ strongly in $H^{1}({\mathbb{R}}^{n})$}. We have also used the hypothesis eq. (99). Thus, for all $\varphi\in C_{0}^{\infty}({\mathbb{R}}^{n})$, eq. (116) implies that $\tilde{w}_{l}$ solves eq. (113) weakly in ${\mathbb{R}}^{n}$.
We next claim that for any $h=1,2,\cdots n$, $\tilde{w}_{l}$ satisfies
(117) $\int_{{\mathbb{R}}^{n}}\left(\nabla\psi^{h}\nabla\tilde{w}_{l}+\psi^{h}\tilde{w}_{l}\right)dz=0$
By direct computation, since $\phi_{j},\psi_{j}\in K^{\perp}_{\epsilon_{j},\bar{\xi}_{j}}$, we get
$|\langle Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}},w_{j}\rangle_{\epsilon_{j}}|=|\langle Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}},\phi_{j}-\psi_{j}-\zeta_{j}\rangle_{\epsilon_{j}}|=|\langle Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}},-\zeta_{j}\rangle_{\epsilon_{j}}|$
[…]
On the other hand
(119) $\langle Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}},w_{j}\rangle_{\epsilon_{j}}=\frac{1}{\epsilon_{j}^{n}}\int_{M}\left[\epsilon_{j}^{2}\nabla_{g}Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h}\nabla_{g}w_{j}+(\epsilon_{j}^{2}\mathbf{c}s_{g}+1)Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h}w_{j}\right]d\mu_{g}$
Meanwhile, using Hölder’s inequality, we compute the second term to be $o(\epsilon_{j})$,
$\left|\frac{1}{\epsilon_{j}^{n}}\int_{M}\epsilon_{j}^{2}\mathbf{c}s_{g}Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h}w_{j}d\mu_{g}\right|\leq c\epsilon_{j}^{2}\left(\frac{1}{\epsilon_{j}^{n}}\int_{M}(Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h})^{2}d\mu_{g}\right)^{1/2}\left(\frac{1}{\epsilon_{j}^{n}}\int_{M}w_{j}^{2}d\mu_{g}\right)^{1/2}$
[…]
since $w_{j}||_{\epsilon_{j}}\rightarrow 1$ as $j\rightarrow\infty$. With this estimation, and rewriting eq. (119) on ${\mathbb{R}}^{n}$:
$\langle Z^{h}_{\epsilon_{j},{\xi_{l}}_{j}},w_{j}\rangle_{\epsilon_{j}}=\frac{1}{\epsilon_{j}^{n}}\int_{M}\left[\epsilon_{j}^{2}\nabla_{g}Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h}\nabla_{g}w_{j}+Z_{\epsilon_{j},{\xi_{l}}_{j}}^{h}w_{j}\right]d\mu_{g}+o(\epsilon_{j})$
$\int_{B(0,r/\epsilon_{j})}\left(\sum_{s,t=1}^{n}g_{{\xi_{l}}_{j}}^{st}(\epsilon_{j}z)\frac{\partial\tilde{w_{l}}_{j}}{\partial z_{s}}\frac{\partial}{\partial z_{t}}(\psi^{h}(z)\chi_{r}(\epsilon_{j}z))+\tilde{w_{l}}_{j}\psi^{h}(z)\chi_{r}(\epsilon_{j}z)\right)|g_{{\xi_{l}}_{j}(\epsilon_{j}z)}|^{1/2}dz+o(\epsilon_{j})$
(120) $\int_{{\mathbb{R}}^{n}}\left(\nabla\psi^{h}\nabla\tilde{w}_{l}+\psi^{h}\tilde{w}_{l}\right)dz+o(1)$
Hence, the claim that $\tilde{w}_{l}$ satisfies (117), follows from (118) and (120). In turn, (113) and (120) imply that $\tilde{w}_{l}=0$.
We now estimate $\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}$ to be $o(1)$ in order to prove (103). We will partition $M$ in different subsets: small neighborhoods of each $\xi_{l}$, $l=1,2,\cdots K$, where we will use $\tilde{w}_{l}=0$, and the complement of these neighborhoods, where we will use the hypothesis (99).
Let $R_{j}=\frac{1}{2}\min\{d_{g}({\xi_{l}}_{j},{\xi_{m}}_{j}),l\neq m\}$. Let $\tilde{M}=\cup_{l=1}^{K}B_{g}({\xi_{l}}_{j},R_{j})$. Hence
(121) $\frac{1}{\epsilon_{j}^{n}}\int_{M}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}=\frac{1}{\epsilon_{j}^{n}}\sum_{l=1}^{K}\int_{B_{g}({\xi_{l}}_{j},R_{j})}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}+\frac{1}{\epsilon_{j}^{n}}\int_{M\setminus\tilde{M}}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}$
For the first term of the right-hand side of (121), for each $l$, since $\tilde{w}_{l}=0$:
$\frac{1}{\epsilon_{j}^{n}}\int_{B_{g}({\xi_{l}}_{j},R_{j})}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}$
$=\int_{B(0,\epsilon_{j}R_{j})}f^{\prime}\left((U(z)+\epsilon_{j}^{2}V(z))\chi_{r}(\epsilon_{j}z)+\sum_{i\neq l}^{K}Y\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z)}{\epsilon_{j}}\right)\chi_{r}(\exp^{-1}_{{\xi_{l}}_{j}}\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))\right)$
$\tilde{w_{l}}_{j}^{2}|g_{{\xi_{l}}_{j}}(\epsilon_{j}z)|^{1/2}dz$
=o(1)=o(1)
Meanwile, for the second term of the right-hand side of (121), by Hölder’s inequality,
$\frac{1}{\epsilon_{j}^{n}}\int_{M\setminus\tilde{M}}f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}})w_{j}^{2}d\mu_{g}$
$\leq\left(\frac{1}{\epsilon_{j}^{n}}\int_{M\setminus\tilde{M}}(f^{\prime}(Y_{\epsilon_{j},\bar{\xi}_{j}}))^{n/2}d\mu_{g}\right)^{\frac{2}{n}}\left(\frac{1}{\epsilon_{j}^{n}}\int_{M\setminus\tilde{M}}w_{j}^{\frac{2n}{n-2}}d\mu_{g}\right)^{\frac{n-2}{n}}$
[…]
$\leq c_{2}\left(\frac{1}{\epsilon_{j}^{n}}\int_{M\setminus\tilde{M}}\left(\sum_{l=1}^{K}U^{p-2}\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}(\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))}{\epsilon_{j}}\right)\chi_{r}^{p-2}\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}(\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))}{\epsilon_{j}}\right)\right)^{n/2}d\mu_{g}\right)^{2/n}+o(\epsilon_{j})$
$\leq c_{2}\frac{1}{\epsilon_{j}^{2}}\sum_{l=1}^{K}\left(\int_{B_{g}({\xi_{l}}_{j},r)\setminus\tilde{M}}U^{\frac{n(p-2)}{2}}\left(\frac{\exp^{-1}_{{\xi_{l}}_{j}}(\exp_{{\xi_{i}}_{j}}(\epsilon_{j}z))}{\epsilon_{j}}\right)d\mu_{g}\right)^{2/n}+o(\epsilon_{j})$
$\leq c_{2}\frac{1}{\epsilon_{j}^{2}}e^{-(p-2)\frac{R_{j}}{\epsilon_{j}}}\sum_{l=1}^{K}\left(\int_{B_{g}({\xi_{l}}_{j},r)\setminus\tilde{M}}d\mu_{g}\right)^{2/n}+o(\epsilon_{j})\leq c_{3}\frac{1}{\epsilon^{2}_{j}}e^{-(p-2)\frac{R_{j}}{\epsilon_{j}}}+o(\epsilon_{j})=o(1)$
where we used the expansion $\left(W_{\epsilon_{j},{\xi_{l}}_{j}}+\epsilon_{j}^{2}V_{\epsilon_{j},{\xi_{l}}_{j}}\right)^{p-2}=W_{\epsilon_{j},{\xi_{l}}_{j}}^{p-2}+o(\epsilon_{j})$, in the third inequality. This proves eq. (103), a contradiction to eq. (102). ∎
We next prove that $R_{\epsilon,\bar{\xi}}$ is suitably bounded.
Lemma 5.2.
There are some $\rho_{0}>0$, $\epsilon_{0}>0$, $C>0$ and $\sigma>0$, such that for any $\rho\in(0,\rho_{0})$, $\epsilon\in(0,\epsilon_{0})$ and $\bar{\xi}\in D^{K_{0}}_{\epsilon,\rho}$, it holds
(122) $R_{\epsilon,\bar{\xi}}||_{\epsilon}\leq\left(C\epsilon^{3}+\sum_{i\neq j}e^{-\frac{(1+\sigma)d_{g}(\xi_{i},\xi_{j})}{2\epsilon}}\right).$
Proof.
We sketch a proof, following [7, Lemma 5.3] and [4, Lemma 3.3]. Let $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})$. Let
(123) $X_{\epsilon,\bar{\xi}}=-\epsilon^{2}\Delta_{g}Y_{\epsilon,\bar{\xi}}+(\epsilon^{2}\mathbf{c}s_{g}+1)Y_{\epsilon,\bar{\xi}}$
this is, $Y_{\epsilon,\bar{\xi}}=i_{\epsilon}^{*}(X_{\epsilon,\bar{\xi}})$.
On the other hand, using eq. (123) and eq. (13), for some $c>0$,
[…]
Hence
(124) $R_{\epsilon,\bar{\xi}}||_{\epsilon}\leq c\left(\left|\left(\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})\right)^{p-1}-\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}\right|_{p^{\prime},\epsilon}+\left|\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\bar{\xi}}\right|_{p^{\prime},\epsilon}\right)$
For the first term of (124), arguing as in Lemma 3.3 in [4], for some $\sigma>0$,
(125) $\left|\left(\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})\right)^{p-1}-\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}\right|_{p^{\prime},\epsilon}=o\left(\sum_{i\neq j}^{K}e^{-\frac{1+\sigma}{2}\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
For the second term of (124), we claim that for some $C>0$,
(126) $\left|\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\bar{\xi}}\right|_{p^{\prime},\epsilon}\leq C\epsilon^{3}$
The estimate of the Lemma follows from (123), (125) and (126).
We next prove (126). We first note that, since $X_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}X_{\epsilon,\xi_{i}}$, we get
$\left|\sum_{i=1}^{K}\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\bar{\xi}}\right|_{p^{\prime},\epsilon}=\left|\sum_{i=1}^{K}\left(\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\xi_{i}}\right)\right|_{p^{\prime},\epsilon}$
(127) $\leq\sum_{i=1}^{K}\left|\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\xi_{i}}\right|_{p^{\prime},\epsilon}$
On the other hand, let $\tilde{X}_{\epsilon,\xi}=X_{\epsilon,\xi}(\exp_{\xi}(z))$. Then, for $z\in B(0,r)\subset{\mathbb{R}}^{n}$,
$\tilde{X}_{\epsilon,\xi}=-\epsilon^{2}\Delta_{g}((U_{\epsilon}+\epsilon^{2}V_{\epsilon})\chi_{r})+(\epsilon^{2}\mathbf{c}s_{g}+1)(U_{\epsilon}\chi_{r}+\epsilon^{2}V_{\epsilon}\chi_{r})$
$U_{\epsilon}^{p-1}(z)\chi_{r}(z)-\epsilon^{2}U_{\epsilon}(z)\Delta\chi_{r}(z)-2\epsilon^{2}\langle\nabla U_{\epsilon}(z),\nabla\chi_{r}(z)\rangle-\epsilon^{4}V_{\epsilon}(z)\Delta\chi_{r}(z)-2\epsilon^{4}\langle\nabla V_{\epsilon}(z),\nabla\chi_{r}(z)\rangle$
$\epsilon^{2}\left[(p-1)|U_{\epsilon}|^{p-2}V_{\epsilon}-\frac{1}{3}R_{ijkl}z_{k}z_{l}\partial_{ij}^{2}U_{\epsilon}+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U_{\epsilon}-\mathbf{c}s_{g}U_{\epsilon}\right]\chi_{r}(z)$
$\epsilon^{2}\left[-\frac{1}{3}R_{ijkl}z_{k}z_{l}\partial_{ij}^{2}U_{\epsilon}+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U_{\epsilon}-\mathbf{c}s_{g}U_{\epsilon}\right]\chi_{r}(z)$
(128) $\epsilon^{4}\frac{1}{3}R_{ijkl}z_{k}z_{l}\partial_{ij}^{2}V_{\epsilon}\chi_{r}(z)+\epsilon^{4}\frac{2}{3}R_{kssj}z_{k}\partial_{j}V_{\epsilon}\chi_{r}(z)+\epsilon^{4}\mathbf{c}s_{g}V_{\epsilon}\chi_{r}+o(\epsilon^{3})$
Then, for each $\xi_{i}$, $i\in i=1,2,\dots,n$,
(129) $\left(\frac{1}{\epsilon^{n}}\int_{M}\left(\left(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}}\right)^{p-1}-X_{\epsilon,\xi_{i}}\right)^{p^{\prime}}d\mu_{g}\right)^{\frac{1}{p^{\prime}}}$
$\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}\left(\left(U_{\epsilon}+\epsilon^{2}V_{\epsilon}\right)^{p-1}\chi^{p-1}-X_{\epsilon,\xi_{i}}(\exp_{\xi_{i}}(z))\right)^{p^{\prime}}|g_{\xi_{i}}(z)|^{\frac{1}{2}}dz\right)^{\frac{1}{p^{\prime}}}$
$\leq C_{0}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}|U_{\epsilon}^{p-1}(z)(\chi_{r}(z)^{p-1}-\chi_{r}(z))|^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}+C_{0}\epsilon^{2}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}U_{\epsilon}^{p^{\prime}}(z)|\Delta\chi_{r}(z)|^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}$
$C_{0}\epsilon^{4}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}V_{\epsilon}^{p^{\prime}}(z)|\Delta\chi_{r}(z)|^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}+C_{0}\epsilon^{2}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}(\nabla U_{\epsilon}(z)\nabla\chi_{r}(z))^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}$
$C_{0}\epsilon^{4}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}(\nabla V_{\epsilon}(z)\nabla\chi_{r}(z))^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}+C_{0}\epsilon^{2}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}|(p-1)U_{\epsilon}^{p-2}(z)V_{\epsilon}(z)(\chi_{r}(z)^{p-1}-\chi_{r}(z))|^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}$
$C_{0}\epsilon^{4}\left(\frac{1}{\epsilon^{n}}\int_{B(0,r)}(\mathbf{c}s_{g}V_{\epsilon}\chi_{r})^{p^{\prime}}dz\right)^{\frac{1}{p^{\prime}}}\leq C\epsilon^{3}$
for some $C>0$. Note that the support of the function $\chi_{r/\epsilon}^{p-1}(y)-\chi_{r/\epsilon}$ and of the derivatives of $\chi_{r/\epsilon}$, are contained in $B(0,\frac{r}{\epsilon})\setminus B(0,\frac{r}{2\epsilon})$. Hence, for some $R>0$, the terms that contain these factors are $o(e^{-R/\epsilon})$, by using the fact that $U$, $V$ and their derivatives decay exponentially.
We are now ready to prove Proposition 3.1.
Proof.
(of Proposition 3.1) To solve eq. (43) we look for a fixed point of the operator $T_{\epsilon,\xi}:H_{\epsilon}\cap K^{\perp}_{\epsilon,\bar{\xi}}\rightarrow H_{\epsilon}\cap K^{\perp}_{\epsilon,\bar{\xi}}$, defined by
$T_{\epsilon,\bar{\xi}}(\phi)=L^{-1}_{\epsilon,\bar{\xi}}\left(N_{\epsilon,\bar{\xi}}(\phi)+R_{\epsilon,\bar{\xi}}\right)$
From lemma 5.1
[…]
Moreover
[…]
By eq. (13)
(132) $N_{\epsilon,\bar{\xi}}(\phi)||_{\epsilon}\leq C|f(Y_{\epsilon,\bar{\xi}}+\phi)-f(Y_{\epsilon,\bar{\xi}})-f^{\prime}(Y_{\epsilon,\bar{\xi}})\phi|_{p^{\prime},\epsilon}$
And by the mean value Theorem there is some $\tau\in(0,1)$ such that if $\phi_{1}||$ and $\phi_{2}||$ are small enough,
$|f(Y_{\epsilon,\bar{\xi}}+\phi_{1})-f(Y_{\epsilon,\bar{\xi}}+\phi_{2})-f^{\prime}(Y_{\epsilon,\bar{\xi}})(\phi_{1}-\phi_{2})|_{p^{\prime},\epsilon}$
$\leq C|f^{\prime}(Y_{\epsilon,\bar{\xi}}+\phi_{2}+\tau(\phi_{1}-\phi_{2}))-f^{\prime}(Y_{\epsilon,\bar{\xi}})(\phi_{1}-\phi_{2})|_{p^{\prime},\epsilon}$
$\leq C|f^{\prime}(Y_{\epsilon,\bar{\xi}}+\phi_{2}+\tau(\phi_{1}-\phi_{2}))-f^{\prime}(Y_{\epsilon,\bar{\xi}})|_{\frac{p}{p-2},\epsilon}\ |(\phi_{1}-\phi_{2})|_{p^{\prime},\epsilon}$
Now, from [4] section 3, eq. (3.34), for $\phi_{1}||$ and $\phi_{2}||$ small enough
$|f^{\prime}(Y_{\epsilon,\bar{\xi}}+\phi_{2}+\tau(\phi_{1}-\phi_{2}))-f^{\prime}(Y_{\epsilon,\bar{\xi}})|_{\frac{p}{p-2},\epsilon}\ |(\phi_{1}-\phi_{2})|_{p^{\prime},\epsilon}\leq|(\phi_{1}-\phi_{2})|_{p^{\prime},\epsilon}$
From eq. (131), this yields
[…]
for $c\in(0,1)$, for $\phi_{1}||$ and $\phi_{2}||$ small enough. Hence, $T_{\epsilon,\bar{\xi}}$ is a contraction map in a small enough ball, centered at 0 in $K^{\perp}_{\epsilon,\bar{\xi}}$.
Moreover, for such fixed point $\phi_{\epsilon,\bar{\xi}}$, from eq.(130)
[…]
This implies the estimate we were looking for
(136) $\phi_{\epsilon,\bar{\xi}}||_{\epsilon}\leq c_{3}\left(\epsilon^{3}+\sum_{i\neq j}e^{-\frac{1+\sigma}{2}\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
Finally, the fact that the map $\bar{\xi}\rightarrow\phi_{\epsilon,\bar{\xi}}$ is $C^{1}$ follows from the Implicit Function Theorem applied to the function
$F(\bar{\xi},\phi)=T_{\epsilon,\bar{\xi}}(\phi)-\phi$
6. Some technical estimates
In this section we provide some estimates that were useful for the last sections. We also prove Proposition 3.3. We begin with a technical estimate that compares the behavior of $\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(w)$ and $\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}$, for small $w$. This is Lemma 5.2 in [4].
Lemma 6.1 (Lemma 5.2 in [4]).
Assume $d_{g}(\xi_{i},\xi_{j})\leq\rho$, $\rho$ small enough. Then
(137) $\tau_{\rho}(w):=\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(w)-\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}},$
$w\in B(0,\rho)$, satisfies,
$\lim_{\rho\rightarrow 0}\sup_{w\in B(0,\rho)\setminus\{0\}}\frac{|\tau_{\rho}(w)|}{w}=1$
In particular, setting $t_{\rho}(z):=\frac{|\tau_{\rho}(\epsilon z)|}{\epsilon}$, for any $\eta>0$, there is some $\rho_{0}>0$, such that for any $\rho\in(0,\rho_{0})$, it holds
(138) $(1-\eta)|z|\leq|t_{\rho}(z)|\leq(1+\eta)|z|,$
for $z\in B(0,\rho/\epsilon)$.
We use the lemma above to provide some useful estimates, based on the exponential decay of $U,V$ and their derivatives.
Lemma 6.2.
If $0<a\leq b$, then for any $\delta$, $0<\delta<1$, provided $\epsilon>0$ and $\rho>0$ are small enough, it holds
$\int_{B(0,\rho/\epsilon)}U^{a}\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)U^{b}(z)dz=o\left(e^{-(a-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right),$
$\int_{B(0,\rho/\epsilon)}V\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)U(z)dz=o\left(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
$\int_{B(0,\rho/\epsilon)}\partial_{k}U(z)\ U\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)dz=o\left(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right),$
$\int_{B(0,\rho/\epsilon)}\partial_{k}U(z)\ V\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)dz=o\left(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right),$
and
$\int_{B(0,\rho/\epsilon)}\partial^{2}_{hk}U(z)\ U\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)dz=o\left(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right).$
Proof.
Recall that $\big|\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}}{\epsilon}\big|=\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}$. For the first estimate, we use the exponential decay of $U(z)$, we have:
$\int_{B(0,\rho/\epsilon)}U^{a}\left(\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\right)U^{b}(z)dz\leq c\int_{{\mathbb{R}}^{n}}e^{-a\big|\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}(\epsilon z)}{\epsilon}\big|-b|z|}dz$
$\leq c\int_{{\mathbb{R}}^{n}}e^{-a\big|\frac{\exp_{\xi_{i}}^{-1}\exp_{\xi_{j}}}{\epsilon}+t_{\rho}(z)\big|-b|z|}dz\leq c\int_{{\mathbb{R}}^{n}}e^{-a\big|\eta+t_{\rho}(z)\big|-b|z|}dz=o\left(e^{-(a-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
where we have used the function $t_{\rho}(z)$, defined in Lemma 6.1, and its propierties (eq. (138)), provided $\rho>0$ is small enough.
The proof of the other estimates is similar, by using the exponential decay of $U$, $V$ and their derivatives. ∎
Lemma 6.3.
It holds
$K_{3}:=\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\ \left[-\epsilon^{2}\Delta W_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})W_{\epsilon,\xi_{i}}-|W_{\epsilon,\xi_{i}}|^{p-1}\right]V_{\epsilon,\xi_{j}}d\mu_{g}=o(\epsilon^{4})$
Proof.
Let $X_{\epsilon,\xi_{i}}:=-\epsilon^{2}\Delta W_{\epsilon,\xi_{i}}+W_{\epsilon,\xi_{i}}$ and $\tilde{X}_{\epsilon,\xi_{i}}:=X_{\epsilon,\xi_{i}}(\exp_{\xi_{i}}(z))$, $z\in B(0,r)$.
Since $\epsilon^{2}\Delta U_{\epsilon}+U_{\epsilon}=U_{\epsilon}^{p-1}$ and by eq. (19) on Lemma 2.1, we get
$\tilde{X}_{\epsilon,\xi_{i}}(z)=U_{\epsilon}^{p-1}(z)\chi_{r}(z)-\epsilon^{2}U_{\epsilon}(z)\Delta\chi_{r}(z)-2\epsilon^{2}\nabla U_{\epsilon}(z)\nabla\chi_{r}(z)$
$\epsilon^{2}(g^{ij}_{\xi_{i}}-\delta_{ij})\partial_{ij}(U_{\epsilon}\chi_{r})-\epsilon^{2}g^{ij}_{\xi_{i}}\Gamma_{ij}^{k}\partial_{k}(U_{\epsilon}\chi_{r})$
Hence, for some $c>0$ and setting $z=\epsilon y$, for $i\neq j$,
$\frac{1}{\epsilon^{n}}\left|\int_{B_{g}(\xi_{i},r)}\left(X_{\epsilon,\xi_{i}}+\epsilon^{2}\mathbf{c}s_{g}W_{\epsilon,\xi_{i}}-W_{\epsilon,\xi_{i}}^{p-1}\right)V_{\epsilon,\xi_{j}}\ d\mu_{g}\right|$
$\frac{1}{\epsilon^{n}}\left|\int_{B(0,r)}\left(\tilde{X}_{\epsilon,\xi_{i}}(z)+\epsilon^{2}\mathbf{c}s_{g}U_{\epsilon}(z)\chi_{r}(z)-U_{\epsilon}^{p-1}(z)\chi_{r}^{p-1}(z)\right)V_{\epsilon,\xi_{j}}(\exp_{\xi_{i}}(z))\ \ |g_{\xi_{i}(z)}|^{\frac{1}{2}}dz\right|$
$\leq\frac{c}{\epsilon^{n}}\int_{B(0,r)}\left|\tilde{X}_{\epsilon,\xi_{i}}(z)+\epsilon^{2}\mathbf{c}s_{g}U_{\epsilon}(z)\chi_{r}(z)-U_{\epsilon}^{p-1}(z)\chi_{r}^{p-1}(z)\right|V_{\epsilon}(\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(z))\ dz$
$c\int_{B(0,r/\epsilon)}\left|\tilde{X}_{\epsilon,\xi_{i}}(\epsilon y)+\epsilon^{2}\mathbf{c}s_{g}U(y)\chi_{r}(\epsilon y)-U^{p-1}(y)\chi_{r}^{p-1}(\epsilon y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\epsilon^{2}\int_{B(0,r/\epsilon)}\left|\mathbf{c}s_{g}U(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|U^{p-1}(y)\left(\chi_{r/\epsilon}^{p-1}(y)-\chi_{r/\epsilon}(y)\right)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}U(y)\left|\Delta\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left\langle\nabla U(y),\nabla\chi_{r/\epsilon}(y)\right\rangle V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(U\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}(U_{\epsilon}\chi_{r})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$H_{0}+H_{1}+H_{2}+H_{3}+H_{4}+H_{5}.$
We claim that each of these terms is $o(\epsilon^{4})$. Note that the support of the function $\chi_{r/\epsilon}^{p-1}(y)-\chi_{r/\epsilon}$ and of the derivatives of $\chi_{r/\epsilon}$, are contained in $B(0,\frac{r}{\epsilon})\setminus B(0,\frac{r}{2\epsilon})$. Hence, for some $R>0$, $H_{1},H_{2},H_{3}=o(e^{-R/\epsilon})$. For $H_{4}$:
$H_{4}=c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(U\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(U(y))\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})U(y)(\partial_{ij}\chi_{r/\epsilon}(y))\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})(\partial_{i}U)(y)(\partial_{j}\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)+o(e^{-R/\epsilon})$
Where we have used $(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})=o(\epsilon^{2}|y|^{2})$ and Lemma 6.2 to argue that the first term is $o(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})$. The other terms involve derivatives of $\chi_{r/\epsilon}$ and hence, as argued above, are $o(e^{-R/\epsilon})$. For $H_{5}$:
$H_{5}=c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}(U\chi_{r\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}U(y)\chi_{r\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)U(y)\partial_{k}(\chi_{r\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)+o(e^{-R/\epsilon})$
Where we have used Lemma 6.2 for the first term, together with the estimate $\Gamma_{ij}^{k}(\epsilon y)=\Gamma_{ij}^{k}(0)+o(\epsilon|y|)$. And we argue, as before, that the second term is $o(e^{-R/\epsilon})$, as it contains derivatives of $\chi_{r/\epsilon}$. Finally, for $H_{0}$,
$H_{0}=c\epsilon^{2}\int_{B(0,r/\epsilon)}\left|\mathbf{c}s_{g}U(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon^{2}\mathbf{c}\int_{B(0,r/\epsilon)}\left|s_{g}U(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy=o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
The Lemma follows from the estimates of $H_{0}$, $H_{1}$, $H_{2}$, $H_{3}$, $H_{4}$ and $H_{5}$.
Lemma 6.4.
It holds
$K_{4}:=\frac{\epsilon^{4}}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}\frac{1}{\epsilon^{n}}\int_{M}\left[-\epsilon^{2}\Delta V_{\epsilon,\xi_{i}}+(1+\epsilon^{2}\mathbf{c}s_{g})V_{\epsilon,\xi_{i}}-(p-1)|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}\right]V_{\epsilon,\xi_{j}}d\mu_{g}=o(\epsilon^{4})$
Proof.
Since $V$ has also exponential decay, the proof is similar to that of Lemma 6.3.
Let $Q_{\epsilon,\xi_{i}}:=-\epsilon^{2}\Delta V_{\epsilon,\xi_{i}}+V_{\epsilon,\xi_{i}}$ and $\tilde{Q}_{\epsilon,\xi_{i}}:=Q_{\epsilon,\xi_{i}}(\exp_{\xi_{i}}(z))$, $z\in B(0,r)$.
Since $\epsilon^{2}\Delta V_{\epsilon}+V_{\epsilon}=(p-1)U_{\epsilon}^{p-2}V_{\epsilon}+R_{kl}z_{k}z_{l}\frac{U_{\epsilon}^{\prime}}{|z|}$, using remark 19 we get
$\tilde{Q}_{\epsilon,\xi_{i}}(z)=-(p-1)U_{\epsilon}^{p-2}(z)V_{\epsilon}(z)\chi_{r}(z)+\epsilon{R_{kl}}_{\xi_{i}}z_{k}z_{l}\frac{U_{\epsilon}^{\prime}}{|z|}\chi_{r}(z)-\epsilon^{2}V_{\epsilon}(z)\Delta\chi_{r}(z)-2\epsilon^{2}\nabla V_{\epsilon}(z)\nabla\chi_{r}(z)$
$\epsilon^{2}(g^{ij}_{\xi_{i}}-\delta_{ij})\partial_{ij}(V_{\epsilon}\chi_{r})(z)-\epsilon^{2}g^{ij}_{\xi_{i}}\Gamma_{ij}^{k}\partial_{k}(V_{\epsilon}\chi_{r})(z)$
Hence, for some $c>0$ and setting $z=\epsilon y$, for $i\neq j$,
$\frac{1}{\epsilon^{n}}\left|\int_{B_{g}(\xi_{i},r)}\left(Q_{\epsilon,\xi_{i}}+\epsilon^{2}\mathbf{c}s_{g}V_{\epsilon,\xi_{i}}-(p-1)|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}\right)V_{\epsilon,\xi_{j}}\ d\mu_{g}\right|$
$\frac{1}{\epsilon^{n}}\left|\int_{B(0,r)}\left(\tilde{Q}_{\epsilon,\xi_{i}}(z)+\epsilon^{2}\mathbf{c}s_{g}V_{\epsilon}(z)\chi_{r}(z)-(p-1)U_{\epsilon}^{p-2}(z)V_{\epsilon}(z)\chi_{r}^{p-1}(z)\right)V_{\epsilon,\xi_{j}}(\exp_{\xi_{i}}(z))\ \ |g_{\xi_{i}(z)}|^{\frac{1}{2}}dz\right|$
$\leq\frac{c}{\epsilon^{n}}\int_{B(0,r)}\left|\tilde{Q}_{\epsilon,\xi_{i}}(z)+\epsilon^{2}\mathbf{c}s_{g}V_{\epsilon}(z)\chi_{r}(z)-(p-1)|U_{\epsilon}|^{p-2}V_{\epsilon}(z)\chi_{r}^{p-1}(z)\right|V_{\epsilon}(\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(z))\ dz$
$c\int_{B(0,r/\epsilon)}\left|\tilde{Q}_{\epsilon,\xi_{i}}(\epsilon y)+\epsilon^{2}\mathbf{c}s_{g}V(y)\chi_{r}(\epsilon y)-(p-1)|U(y)|^{p-2}V(y)\chi_{r}^{p-1}(\epsilon y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\int_{B(0,r/\epsilon)}\left|U^{p-2}(y)V(y)\left(\chi_{r/\epsilon}^{p-1}(y)-\chi_{r/\epsilon}(y)\right)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}V(y)\left|\Delta\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left\langle\nabla V(y),\nabla\chi_{r/\epsilon}(y)\right\rangle V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(V\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}(V_{\epsilon}\chi_{r})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon\int_{B(0,r/\epsilon)}R_{kl}y_{k}y_{l}\frac{U^{\prime}(y)}{|y|}V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon^{2}\int_{B(0,r/\epsilon)}\left|\mathbf{c}s_{g}V(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$L_{1}+L_{2}+L_{3}+L_{4}+L_{5}+L_{6}+L_{7}.$
Note that the support of the function $\chi_{r/\epsilon}^{p-1}(y)-\chi_{r/\epsilon}$ and of the derivatives of $\chi_{r/\epsilon}$, are contained in $B(0,\frac{r}{\epsilon})\setminus B(0,\frac{r}{2\epsilon})$. Hence, for some $R>0$,
(139) $L_{1},L_{2},L_{3}=o(e^{-R/\epsilon})$
For $L_{4}$:
$L_{4}=c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(V\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})\partial_{ij}(V(y))\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})V(y)(\partial_{ij}\chi_{r/\epsilon}(y))\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\int_{B(0,r/\epsilon)}\left|(g^{ij}_{\xi_{i}}(\epsilon y)-\delta_{ij})(\partial_{i}V)(y)(\partial_{j}\chi_{r/\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)+o(e^{-R/\epsilon})$
Where we have used Lemma 6.2 (to argue that the first term is $o(e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}})$). And we argue as before, that the remainder of the terms are $o(e^{-R/\epsilon})$, since they contain derivatives of $\chi_{r/\epsilon}$. For $L_{5}$:
$L_{5}=c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}(V\chi_{r\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$\leq c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)\partial_{k}(V)(y)\chi_{r\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon\int_{B(0,r/\epsilon)}\left|g^{ij}_{\xi_{i}}(\epsilon y)\Gamma_{ij}^{k}(\epsilon y)V(y)\partial_{k}(\chi_{r\epsilon})(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)+o(e^{-R/\epsilon})$
Where we have used Lemma 6.2 for the first term, together with the estimate $\Gamma_{ij}^{k}(\epsilon y)=\Gamma_{ij}^{k}(0)+o(\epsilon|y|)$. And we argue, as before, that the second term is $o(e^{-R/\epsilon})$, as it contains derivatives of $\chi_{r/\epsilon}$. For $L_{6}$:
$L_{6}=c\epsilon\int_{B(0,r/\epsilon)}R_{kl}y_{k}y_{l}\frac{U^{\prime}(y)}{|y|}V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy=o\left(\epsilon e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
by Lemma 6.2. For $L_{7}$:
$L_{7}=c\epsilon^{2}\int_{B(0,r/\epsilon)}\left|\mathbf{c}s_{g}V(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy$
$c\epsilon^{2}\mathbf{c}\int_{B(0,r/\epsilon)}\left|s_{g}V(y)\chi_{r/\epsilon}(y)\right|V\left(\frac{\exp_{\xi_{j}}^{-1}\exp_{\xi_{i}}(\epsilon y)}{\epsilon}\right)\ dy=o\left(\epsilon^{2}e^{-(1-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
Where we have used Lemma 6.2.
The Lemma follows from the estimates of $L_{0}$, $L_{1}$, $L_{2}$, $L_{3}$, $L_{4}$, $L_{5}$ and $L_{6}$.
The following estimate will be useful.
Lemma 6.5.
Let $\bar{\xi}=(\xi_{1},\xi_{2},\dots,\xi_{n})\in D_{\epsilon,\rho}^{K}$. For $q>0$, it holds
$\frac{1}{\epsilon^{n}}\int_{M}\left(|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)\ d\mu_{g}=o(\epsilon^{2})$
provided $\rho>0$ is small enough.
Proof.
We follow the ideas of (the proof of) Lemma 3.3 in [4] and some estimates in (the proof of) Lemma 4.1 in the same article.
For $0<q\leq 2$, we use the following inequality. For any $a,b>0$, there is some $c>0$ such that
(140) $\left||a+b|^{q}-a^{q}-b^{q}\right|\leq c\ a^{\frac{q}{2}}b^{\frac{q}{2}}$
Hence, applying (140) repeatedly, we have,
$\frac{1}{\epsilon^{n}}\left(\int_{M}|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)d\mu_{g}\leq c\frac{1}{\epsilon^{n}}\sum_{i=1}^{K}\sum_{j=i+1}^{K}\int_{M}W_{\epsilon,\xi_{i}}^{\frac{q}{2}}W_{\epsilon,\xi_{j}}^{\frac{q}{2}}d\mu_{g}$
Moreover, for $i\neq j$,
$\frac{1}{\epsilon^{n}}\int_{M}W_{\epsilon,\xi_{i}}^{\frac{q}{2}}W_{\epsilon,\xi_{j}}^{\frac{q}{2}}d\mu_{g}=\int_{B(0,r/\epsilon)}U^{\frac{q}{2}}(z)\chi_{r}^{\frac{q}{2}}(z)U^{\frac{q}{2}}\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)\chi_{r}^{\frac{q}{2}}(\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z)))|g_{\xi_{i}}(\epsilon z)|^{\frac{1}{2}}dz$
$\leq c\int_{{\mathbb{R}}^{n}}U^{\frac{q}{2}}(z)U^{\frac{q}{2}}\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)dz=o\left(\sum_{i\neq j}^{K}e^{-(\frac{q}{2}-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
where the last inequality follows from Lemma 6.2. For $q>2$, we use the following inequalities. For any $a,b>0$, there is some $c>0$ such that
(141) $\left||a+b|^{q}-a^{q}-b^{q}-pa^{q-1}b-qab^{q-1}\right|\leq\begin{cases}c\ a^{\frac{q}{2}}b^{\frac{q}{2}}\ \ \text{if}\ 2<q\leq 3\\ c(a^{q-2}b^{2}+a^{2}b^{q-2})\ \ \text{if}\ \ q>3\end{cases}$
Note also that, from Lemma 6.2, for $q>2$,
(142) $\frac{1}{\epsilon^{n}}\sum_{i\neq j}^{K}\int_{M}W_{\epsilon,\xi_{i}}^{\frac{q}{2}}W_{\epsilon,\xi_{j}}^{\frac{q}{2}}d\mu_{g}=o\left(\sum_{i\neq j}^{K}e^{-(\frac{q}{2}-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)=o\left(\sum_{i\neq j}^{K}e^{-(1+\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
(143) $\frac{1}{\epsilon^{n}}\sum_{i\neq j}^{K}\int_{M}W_{\epsilon,\xi_{i}}^{q-2}W_{\epsilon,\xi_{j}}^{2}d\mu_{g}=o\left(\sum_{i\neq j}^{K}e^{-(\min\{q-2,2\}-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)=o\left(\sum_{i\neq j}^{K}e^{-(1+\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
Hence, using eq. (140) with $a=|\sum_{i=2}^{K}W_{\epsilon,\xi_{i}}|$ and $b=W_{\epsilon,\xi_{1}}$, and then, $a=|\sum_{i=j+1}^{K}W_{\epsilon,\xi_{i}}|$ and $b=W_{\epsilon,\xi_{j}}$, in a successive manner, we have,
$\frac{1}{\epsilon^{n}}\left(\int_{M}|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)d\mu_{g}$
$\frac{1}{\epsilon^{n}}\left(\int_{M}|\sum_{i=2}^{K}W_{\epsilon,\xi_{i}}|^{q}-\sum_{i=2}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)d\mu_{g}+q\frac{1}{\epsilon^{n}}\int_{M}|\sum_{i=2}^{K}W_{\epsilon,\xi_{i}}|^{q-1}W_{\epsilon,\xi_{1}}d\mu_{g}+q\frac{1}{\epsilon^{n}}\int_{M}W_{\epsilon,\xi_{1}}^{q-1}\sum_{i=2}^{K}W_{\epsilon,\xi_{i}}d\mu_{g}$
$o\left(\sum_{i\neq j}^{K}e^{-(1+\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
$q\frac{1}{\epsilon^{n}}\int_{M}\sum_{j=1}^{K-1}|\sum_{i=j+1}^{K}W_{\epsilon,\xi_{i}}|^{q-1}W_{\epsilon,\xi_{j}}d\mu_{g}+q\frac{1}{\epsilon^{n}}\int_{M}\sum_{i<j}^{K}W_{\epsilon,\xi_{i}}^{q-1}W_{\epsilon,\xi_{j}}d\mu_{g}$
$o\left(\sum_{i\neq j}^{K}e^{-(1+\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
where we have used eq. (142) if $2<q\leq 3$ and (143) if $q>3$. Now, from eq. (4.8) in [4], for $q>2$,
(144) $\frac{1}{\epsilon^{n}}\int_{M}\sum_{j=1}^{K-1}|\sum_{i=j+1}^{K}W_{\epsilon,\xi_{i}}|^{q-1}W_{\epsilon,\xi_{j}}d\mu_{g}=\sum_{i=j+1}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}\xi_{j}}{\epsilon}\right)+o(\epsilon^{2})$
Moreover, being $\bar{\xi}\in D^{K}_{\epsilon,\rho}$, then $\sum_{i\neq}U\left(\frac{\exp_{\xi_{i}}^{-1}\xi_{j}}{\epsilon}\right)<\epsilon^{4}$. Hence, since $\gamma_{ij}$ is bounded, it follows that (144) is at least $o(\epsilon^{2})$. On the other hand,
$\frac{1}{\epsilon^{n}}\sum_{i<j}^{K}\int_{M}W_{\epsilon,\xi_{i}}^{q-1}W_{\epsilon,\xi_{j}}d\mu_{g}=o\left(\sum_{i<j}^{K}e^{-((q-1)-\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)=o\left(\sum_{i<j}^{K}e^{-(1+\delta)\frac{d_{g}(\xi_{i},\xi_{j})}{\epsilon}}\right)$
Lemma 6.6.
It holds
$K_{5}:=\frac{\epsilon^{2}}{\epsilon^{n}}\left[\sum_{j=1}^{K}\int_{M}-|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}+\sum_{i=1}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{i}}\ d\mu_{g}+\sum_{i\neq j}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}\right]=o(\epsilon^{4})$
Proof.
Note that $K_{5}$ is equivalent to,
$\epsilon^{2}\frac{1}{\epsilon^{n}}\sum_{j=1}^{K}\left[\int_{M}-|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}d\mu_{g}+\int_{M}\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{p-1}V_{\epsilon,\xi_{j}}\ d\mu_{g}\right]$
So that to prove the Lemma, it suffices to prove that for any $j\in\{1,2,\dots,K\}$,
$\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)V_{\epsilon,\xi_{j}}\ d\mu_{g}=o(\epsilon^{2})$
with q=p1q=p-1. Since $V_{\epsilon,\xi_{j}}$ is bounded for each $j\in\{1,2,\dots,K\}$, and by Lemma 6.5, using q=p1>0q=p-1>0:
$\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)V_{\epsilon,\xi_{j}}\ d\mu_{g}$
$\leq c\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)\ d\mu_{g}=o(\epsilon^{2})$
Lemma 6.7.
It holds, $K_{6}=o(\epsilon^{4})$, with
$K_{6}:=$
$\frac{\epsilon^{4}}{2}\frac{(p-1)}{\epsilon^{n}}\left[\sum_{l=1}^{K}\sum_{j=1}^{K}\int_{M}-|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}+\sum_{i=1}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}+\sum_{i\neq j}^{K}\int_{M}|W_{\epsilon,\xi_{i}}|^{p-2}V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}\right]$
Proof.
Note that $K_{6}$ is equivalent to,
$\frac{\epsilon^{4}}{2}(p-1)\frac{1}{\epsilon^{n}}\int_{M}-\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{p-2}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{p-2}\right)V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}$
Hence, it suffices to prove that for any $j,l\in\{1,2,\dots,K\}$,
$\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}=o(1)$
with q=p2q=p-2. Since $V_{\epsilon,\xi_{j}}$ is uniformily bounded for each $j,l\in\{1,2,\dots,K\}$, and by Lemma 6.5, using q=p2>0q=p-2>0:
$\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)V_{\epsilon,\xi_{j}}V_{\epsilon,\xi_{l}}d\mu_{g}\leq c\frac{1}{\epsilon^{n}}\int_{M}\left(\left|\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}\right|^{q}-\sum_{i=1}^{K}|W_{\epsilon,\xi_{i}}|^{q}\right)d\mu_{g}=o(1)$
The following are estimations of the functions $Z^{l}_{\epsilon,\xi}$, defined in eq. (10), they were used to prove Lemma 5.1 and were proved in [4].
Lemma 6.8 (Lemma 5.3 in [4]).
It holds
(145) $\frac{\partial}{\partial y_{h}^{j}}Z^{l}_{\epsilon,\xi_{i}(y_{i})}=\frac{\partial}{\partial y_{h}^{j}}W_{\epsilon,\xi_{i}(y^{i})}=0,\text{ if }i\neq j$
(146) $\left\langle Z^{l}_{\epsilon,\xi_{j}},\frac{\partial}{\partial y_{h}^{i}}W_{\epsilon,\xi_{i}(y^{i})}\right\rangle_{\epsilon}=o(1),\text{ if }i\neq j$
(147) $\frac{\partial}{\partial y_{h}^{i}}Z^{l}_{\epsilon,\xi_{i}(y_{i})}||_{\epsilon}=o\left(\frac{1}{\epsilon}\right),$
(148) $\left\langle Z^{l}_{\epsilon,\xi_{i}},Z^{h}_{\epsilon,\xi_{j}}\right\rangle_{\epsilon}=o(1),\text{ if }i\neq j$
(149) $\left\langle Z^{l}_{\epsilon,\xi_{i}},Z^{h}_{\epsilon,\xi_{i}}\right\rangle_{\epsilon}=c\delta_{lh}+o(1),$
where $c=\int_{{\mathbb{R}}^{n}}\left(|\nabla\psi^{l}|^{2}+(\psi^{l})^{2}\right)dz$ is a positive constant.
Finally, we sketch a proof of Proposition 3.3. We argue as in Lemma 4.1 in [12] and [19], Proposition 3.1.
Proof.
(Of Proposition 3.3)
Let $x^{\alpha}\in B(0,r)$, for $\alpha=1,\dots K$ and let $y^{\alpha}=\exp_{\xi_{\alpha}}(x^{\alpha})$, $\bar{y}=(y^{1},y^{2},\dots,y^{K})\in M^{K}$. Since $\bar{\xi}$ is a critical point of $\bar{J}_{\epsilon}$, for $\alpha=1,\dots,K$, $i=1,\dots,n$,
(150) $\frac{\partial}{\partial x_{i}^{\alpha}}\bar{J}_{\epsilon}(\bar{y}(x))=0$
$S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})=\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\}+\Pi_{\epsilon,\xi}\left\{S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\}$
Note that $\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\}=0$ by construction of $\phi_{\epsilon,\bar{y}(x)}$. Meanwhile
$\Pi_{\epsilon,\xi}\left\{S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\}=\sum_{i,\alpha}C_{\epsilon}^{i,\alpha}Z^{i}_{\epsilon,y^{\alpha}}$
for some functions $C_{\epsilon}^{i,\alpha}:B(0,r)^{K}\rightarrow{\mathbb{R}}$. We now prove that $C_{\epsilon}^{i,\alpha}(0)=0$ for each $i,\alpha$. Note that for each $i,\alpha$,
$0=\frac{\partial}{\partial x_{i}^{\alpha}}\bar{J}_{\epsilon}(\bar{y}(x))=\bar{J}_{\epsilon}^{\prime}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\left[\frac{\partial}{\partial x_{i}^{\alpha}}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right]$
$\left\langle S_{\epsilon}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)}),\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\rangle_{\epsilon}$
(151) $\left\langle\sum_{k,\beta}C_{\epsilon}^{k,\beta}(0)Z^{k}_{\epsilon,y^{\beta}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}(Y_{\epsilon,\bar{y}(x)}+\phi_{\epsilon,\bar{y}(x)})\right\rangle_{\epsilon}$
Now, since $\phi_{\epsilon,\bar{\xi}(y)}\in K^{\perp}_{\epsilon,\bar{\xi}(y)}$, hence $\left\langle Z^{k}_{\epsilon,y^{\beta}},\phi_{\epsilon,\bar{y}(x)}\right\rangle=0$. Then
(152) $\liminf_{\epsilon\rightarrow 0}\Big|\left\langle Z^{k}_{\epsilon,y^{\beta}},\left(\frac{\partial}{\partial x_{i}^{\alpha}}\phi_{\epsilon,\bar{y}(x)}\right)\Big|_{y=0}\right\rangle_{\epsilon}\Big|=\liminf_{\epsilon\rightarrow 0}\Big|-\left\langle\frac{\partial}{\partial x_{i}^{\alpha}}Z^{k}_{\epsilon,y^{\beta}}\Big|_{y=0},\phi_{\epsilon,\bar{y}(x)}\right\rangle_{\epsilon}\Big|$
[…]
where the last inequality follows from eq. (45) in Proposition 3.1. Now,
$\left\langle\sum_{k,\beta}C_{\epsilon}^{k,\beta}(0)Z^{k}_{\epsilon,y^{\beta}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,\bar{y}(x)}\right\rangle_{\epsilon}=\left\langle\sum_{k,\beta}C_{\epsilon}^{k,\beta}(0)Z^{k}_{\epsilon,y^{\beta}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}$
$\left\langle\sum_{k}C_{\epsilon}^{k,\alpha}(0)Z^{k}_{\epsilon,y^{\alpha}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}+\left\langle\sum_{k,\beta\neq\alpha}C_{\epsilon}^{k,\beta}(0)Z^{k}_{\epsilon,y^{\beta}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}$
Note that for $\alpha\neq\beta$, by the exponential decay of $U$ and $V$,
$\lim_{\epsilon\rightarrow 0}\left\langle\sum_{k,\beta\neq\alpha}C_{\epsilon}^{k,\beta}(0)Z^{k}_{\epsilon,y^{\beta}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}=0$
On the other hand,
$\left\langle\sum_{k}C_{\epsilon}^{k,\alpha}(0)Z^{k}_{\epsilon,y^{\alpha}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}$
$\left\langle C_{\epsilon}^{i,\alpha}(0)Z^{i}_{\epsilon,y^{\alpha}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}+\left\langle\sum_{k\neq i,\alpha}C_{\epsilon}^{k,\alpha}(0)Z^{k}_{\epsilon,y^{\alpha}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}$
And since $\lim_{\epsilon\rightarrow 0}\epsilon\langle Z^{i}_{\epsilon,\xi},\frac{\partial Y_{\epsilon,\xi}}{\partial x^{k}}\rangle_{\epsilon}=\delta_{ik}C$, then
$\lim_{\epsilon\rightarrow 0}\epsilon\left\langle\sum_{k}C_{\epsilon}^{k,\alpha}(0)Z^{k}_{\epsilon,y^{\alpha}},\frac{\partial}{\partial x_{i}^{\alpha}}\Big|_{x=0}Y_{\epsilon,y^{\alpha}(x)}\right\rangle_{\epsilon}=C_{\epsilon}^{i,\alpha}(0)C$
It follows from (151) and (152) that $C_{\epsilon}^{i,\alpha}(0)=0$.